### v2

# Asymptotic Behavior of Integral Projection Models via Genealogical Quantities

Abstract.
Multi-state structured population models, including integral projection models (IPMs) and age-structured McKendrick equations, link individual life histories to population growth and composition, yet the demographic meaning of their dominant eigenstructure can be difficult to interpret. A main goal of this paper is to derive interpretable demographic indicators for multi-state heterogeneity—in particular expected generation numbers, which act as an effective genealogical memory length (in generations) of the ancestry-weighted contributions driving growth—together with type reproduction numbers and generation intervals, directly from life-history transition kernels.
To this end we develop a determinant-free genealogical framework based on a reference-point operator, a rank-one construction at the kernel level that singles out a biologically chosen reference state and organizes lineages by their contributions relative to that state. This yields stable distributions and reproductive values as convergent series of iterated kernels, and leads to an Euler–Lotka-like characteristic equation expressed by reference-point moments. The resulting expansion admits a closed combinatorial form via ordinary partial Bell polynomials, providing a direct bridge from transition kernels to genealogical quantities.
We extend the approach to multi-state McKendrick equations and show how these indicators quantify how population scale and composition are determined by ancestry-weighted initial-state information. The framework avoids restrictive Hilbert–Schmidt assumptions and clarifies how temporal memory and multi-type heterogeneity emerge from cross-generational accumulation, yielding a unified and interpretable route from transition kernels to multi-state demographic indicators.
International Relations, National Institute of Population and Social Security Research, 2-2-3 Uchisaiwai-cho, Chiyoda-ku, Tokyo 100-0011, Japan
Academic Support Center, Kogakuin University, 2665-1 Nakano-machi, Hachioji-shi, Tokyo 192-0015, Japan
Department of Applied Mathematics, National Yang Ming Chiao Tung University, Hsinchu 30010, Taiwan (R.O.C.)
1. Introduction
The Integral Projection Model (IPM) is a mathematical framework for describing transitions in populations with continuous state variables (Ellner and Rees, 2006). In discrete-time settings, IPMs have been developed and applied widely in ecology and demography. For example, Coulson et al. (2010) studied trait evolution, while Coulson and Tuljapurkar (2008) analyzed asymptotic population dynamics under environmental variability. These and other studies (White et al., 2016; Nicole et al., 2011; Doak et al., 2021; Merow et al., 2014; Elderd and Miller, 2016) have linked spectral theory to equilibrium states, generation times, reproductive values, and related quantities, often supported by empirical analyses using plant and animal data. A continuous-time analogue, the Path Integral Model (Oizumi and Takada, 2013; Oizumi, 2022b), derives population-level dynamics from stochastic models of individual life histories and can incorporate optimization principles such as $r/K$-selection (Oizumi et al., 2016). It can be viewed as a variant of the multi-state McKendrick equation, which generalizes the McKendrick–von Foerster PDE by including state variables beyond age.
From a biological and demographic perspective, the dominant eigenstructure of an IPM is valuable not only because it determines the long-run growth rate and stable structure, but because it provides a principled way to connect initial states (traits, stages, or age–state conditions) to population-scale outcomes through the accumulation of contributions across generations. In multi-state McKendrick-type models, this viewpoint naturally emphasizes genealogical information: how lineages initiated from a given age–state condition propagate, reproduce, and redistribute mass through the life-cycle, thereby shaping the stable age–state distribution and the reproductive value.
Mathematically, these quantities are governed by the spectral radius of the associated positive integral operator and its eigenfunctions, which are often characterized by a Fredholm integral equation of the second kind. Fredholm’s original solution (Fredholm, 1903), expressed via the Fredholm determinant, is foundational in functional analysis, but determinant-based formulations can be computationally expensive and may provide limited biological transparency in applied settings. For practical inference, many studies adopt quadrature-based matrix approximations (Ellner and Rees, 2006; White et al., 2016) that discretize the state space; however, discretization requires care because it may not preserve qualitative properties of the original continuous model.
In demographic modeling, a classical alternative to determinant calculations is the Euler–Lotka equation (Euler, 1760; Sharpe and Lotka, 1911), which is equivalent to the characteristic equation for the intrinsic growth rate but is more tractable and biologically transparent, relating directly to net reproduction and generation-time distributions (Inaba, 2017). Recently, Oizumi et al. (2022a) extended a classical Markov chain idea (Chung, 1960) to irreducible nonnegative matrices, expressing eigenvectors associated with the Frobenius root explicitly in terms of matrix entries and taboo probabilities.
Here, we use the term “taboo” in precisely this Markov-chain sense: in a discrete-state chain, taboo probabilities describe transitions killed upon hitting a designated state and lead to clean decompositions into first-visit and renewal contributions. In continuous-state IPMs, however, a literal taboo at a single point cannot be imposed purely at the level of $L^{1}$-type kernel spaces, because point evaluation is not automatically a bounded functional. In this paper we therefore continuousize the Markov-chain analogy by introducing a reference-point construction at the kernel level that makes point evaluation well posed within the biological kernel setting. We postpone a detailed discussion of the structural analogy with taboo probabilities and its continuous-state interpretation to Section 3.
Aim and contribution. Motivated by this determinant-free and interpretation-oriented viewpoint, the aim of this paper is to develop an explicit, determinant-free formulation of eigenfunctions for positive-kernel Fredholm equations that arise in continuous-state IPMs, and to connect the resulting representations to genealogical and demographic quantities in multi-state McKendrick models.
Our key device is a reference-point operator (a rank-one operator at the kernel level) constructed from a fixed reference pair $(x_{0},y_{0})$. This operator algebraically implements a Markov-chain–inspired “taboo” decomposition at the reference point in a way that remains well posed for continuous-state kernels, and it leads to a family of reference-point iterates $\Gamma_{n}$ generated by $(\mathbf{K}-\mathbf{P})$. The resulting Neumann-series/resolvent expansion replaces determinant expressions by sums over iterated kernels, so each term admits a direct genealogical reading as a multi-generation contribution. Moreover, the combinatorics of the iterates admit a closed representation in terms of ordinary partial Bell polynomials, yielding an explicit expression of $\Gamma_{n}$ and the associated reference-point moments. These moments in turn provide a characteristic equation in an Euler–Lotka-like form, which is well suited to positivity-based analysis and biological interpretation.
Scope of applications. Applied to a simple discrete-time IPM, the formulation gives explicit representations of the stable state distribution and reproductive value. Applied to multi-state McKendrick equations, it yields genealogical interpretations of key demographic quantities (e.g., stable age–state distributions and reproductive values) and enables the direct computation of demographic indicators that are driven by ancestry and initial-state information—such as type reproduction numbers, expected generation numbers, and generation intervals—directly from continuous-state models, without passing through discretization.
Organization. The remainder of this paper is organized as follows. In Section 2, we derive a solution to the Fredholm integral equation with a positive integral kernel without relying on the Fredholm determinant, based on the reference-point operator and the $(\mathbf{K}-\mathbf{P})$-iterates. Section 3 applies the theory to a simple IPM and discusses general properties of its eigenstructure and biological interpretation, including the Markov-chain taboo analogy and its continuous-state implementation. Section 4 applies the results to the multi-state McKendrick equation and develops demographic indicators arising from the genealogical expansion. Section 5 summarizes the results and discusses previous perspectives on IPMs as well as new challenges for future research.
2. Reconstruction of the solution to the nonnegative Fredholm integral equation
In this chapter, we consider the Fredholm integral equation with a nonnegative kernel, a formulation frequently encountered in theoretical studies of structured population dynamics.
2.1. Kernel space, kernels, and iterates
We first specify the kernel space on which the reference-point construction is performed. Let $\Omega^{d}\subseteq\mathbb{R}^{d}$ be a measurable domain and write $\mu$ for the Lebesgue measure on $\Omega^{d}$. We treat kernels as measurable functions on $\Omega^{d}\times\Omega^{d}$ and use the mixed-norm space
(1) $\mathcal{X}:=L^{\infty}\!\big(\Omega^{d}_{y};L^{1}(\Omega^{d}_{x})\big),\qquad\|F\|_{\mathcal{X}}:=\operatorname*{ess\,sup}_{y\in\Omega^{d}}\int_{\Omega^{d}}|F(x,y)|\,dx.$
Thus $F\in\mathcal{X}$ means that $x\mapsto F(x,y)$ is integrable for a.e. $y$, with an $L^{1}$-bound uniform in $y$.
Definition 2.1 (Admissible kernels).
Let $\mathbb{K}$ be the class of kernels $K:\Omega^{d}\times\Omega^{d}\to\mathbb{R}$ such that:
  • Positivity: $K(x,y)>0$ for all $x,y\in\Omega^{d}$.
  • Mixed-norm integrability: $K\in\mathcal{X}$, i.e.
    $\operatorname*{ess\,sup}_{y\in\Omega^{d}}\int_{\Omega^{d}}|K(x,y)|\,dx<\infty.$
  • Uniform $L^{\infty}$-boundedness in the second variable: there exists $M>0$ such that
    $\operatorname*{ess\,sup}_{x\in\Omega^{d}}\|K(x,\cdot)\|_{L^{\infty}(\Omega^{d})}\leq M.$
  • Continuity: $K$ is (jointly) continuous on $\Omega^{d}\times\Omega^{d}$.
Integral operator and iterated kernels.
Given $K\in\mathbb{K}$, define the integral operator $\mathbf{K}$ acting on the first variable by
(2) $(\mathbf{K}f)(x):=\int_{\Omega^{d}}K(x,\xi)\,f(\xi)\,d\xi,\qquad f\in L^{1}(\Omega^{d}).$
For later use, we also let $\mathbf{K}$ act on kernels $F\in\mathcal{X}$ by
(3) $(\mathbf{K}F)(x,y):=\int_{\Omega^{d}}K(x,\xi)\,F(\xi,y)\,d\xi,\qquad F\in\mathcal{X},$
whenever the integral is finite. (In particular, for $F\in\mathcal{X}$ and $K\in\mathbb{K}$, the right-hand side is well-defined for a.e. $y$.)
Point evaluation as a focal-state taboo (biological motivation and well-posedness).
Fix a focal state $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$. In many IPM applications this represents a sentinel trait/stage (threshold, target state, or a designated reference state) and we want to decompose genealogical paths by whether they pass through $(x_{0},y_{0})$. To encode a literal taboo at a single state, we need point evaluation $F\mapsto F(x_{0},y_{0})$ to be a bounded functional on the kernel space.
Choose an open neighborhood $U\subset\Omega^{d}\times\Omega^{d}$ of $(x_{0},y_{0})$ with $|U|<\infty$ and define
$\|F\|_{\infty,U}:=\sup_{(x,y)\in U}|F(x,y)|,\qquad\mathcal{X}_{\star}:=\{F\in\mathcal{X}:\ \|F\|_{\infty,U}<\infty\},\qquad\|F\|_{\mathcal{X}_{\star}}:=\|F\|_{\mathcal{X}}+\|F\|_{\infty,U}.$
Then point evaluation at the focal state is bounded on $\mathcal{X}_{\star}$:
$|F(x_{0},y_{0})|\leq\|F\|_{\infty,U}\leq\|F\|_{\mathcal{X}_{\star}}.$
Remark (terminology and forward reference). We call this construction a taboo at the focal state $(x_{0},y_{0})$ in analogy with taboo probabilities (killed/avoided paths) in Markov and branching models; the precise structural correspondence and its demographic interpretation are discussed in Section 3.
Rank-one point-evaluation correction and a taboo operator on kernels.
For $F\in\mathcal{X}_{\star}$, define the rank-one correction operator $\mathbf{P}:\mathcal{X}_{\star}\to\mathcal{X}_{\star}$ by
$(\mathbf{P}F)(x,y):=K(x,y)\,F(x_{0},y_{0}).$
Using the already-defined action of $\mathbf{K}$ on kernels (3), define the taboo operator $\mathbf{A}:\mathcal{X}_{\star}\to\mathcal{X}_{\star}$ by
$\mathbf{A}:=\mathbf{K}-\mathbf{P},\qquad\text{i.e.}\qquad(\mathbf{A}F)(x,y)=\int_{\Omega^{d}}K(x,\xi)\,F(\xi,y)\,d\xi-K(x,y)\,F(x_{0},y_{0}).$
Biologically, $\mathbf{P}$ extracts the genealogical mass that visits the focal state $(x_{0},y_{0})$ and redistributes it via the same one-step kernel $K$, so $\mathbf{A}$ updates genealogical kernels with a literal taboo at the focal state.
Standing notation in Chapter 2.
Throughout Chapter 2, we fix an oprator norm:
$N_{0}:=\|\mathbf{K}\|_{op}.$
Whenever a reference pair $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$ is fixed, we write $K_{0}:=K(x_{0},y_{0})$.
We define the iterated kernels by the convention
(4) $K^{(1)}(x,y):=K(x,y),\qquad K^{(n+1)}(x,y):=(\mathbf{K}K^{(n)})(x,y)=\int_{\Omega^{d}}K(x,\xi)\,K^{(n)}(\xi,y)\,d\xi,\quad n\geq 1.$
Equivalently, $K^{(n)}=\mathbf{K}^{\,n-1}K$ for $n\geq 1$.
Remark 2.2 (Point evaluation).
Although point evaluation is not a bounded functional on $L^{1}(\Omega^{d})$, it is well-defined for kernels as functions. Since $K$ is continuous, $K(x_{0},y_{0})$ is meaningful for every $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$. In the reference-point construction below, evaluation at $(x_{0},y_{0})$ is used only at the kernel level to define the rank-one operator $\mathbf{P}$ generating the taboo-type iterates.
Remark 2.3.
The class $\mathbb{K}$ is designed for IPMs in mathematical biology and does not assume Hilbert–Schmidt and trace class structure. For comparison, many classical Fredholm-determinant arguments are most transparent in the subclass
$\mathbb{K}_{2}:=\left\{K\in\mathbb{K}\,\middle|\,\int_{\Omega^{d}\times\Omega^{d}}|K(x,y)|^{2}\,dx\,dy<\infty\,and\int_{\Omega^{d}}|K(x,x)|^{2}\,dx<\infty\right\},$
whose members define Hilbert–Schmidt operators on $L^{2}(\Omega^{d})$. Our main construction below does not require $K\in\mathbb{K}_{2}$.
We consider the eigenvalue equation
(5) $w(x,y;\lambda_{0})=\frac{1}{\lambda_{0}}\int_{\Omega^{d}}K(x,\xi)\,w(\xi,y;\lambda_{0})\,d\xi=\frac{1}{\lambda_{0}}(\mathbf{K}w)(x,y;\lambda_{0}).$
Definition 2.4 (Partial Bell polynomials).
Let $Z_{1},Z_{2},\dots$ be a sequence of variables. For integers $\mu\geq 0$ and $\nu\geq 0$, the ordinary partial Bell polynomial $\widehat{B}_{\mu,\nu}(Z_{1},\dots,Z_{\mu};x,y)$ is defined by
$\widehat{B}_{\mu,\nu}(Z_{1},\dots,Z_{\mu};x,y):=\sum_{\begin{subarray}{c}(m_{j})_{1\leq j\leq\mu}\geq 0\\ \sum_{j=1}^{\mu}m_{j}=\nu,\ \sum_{j=1}^{\mu}j\,m_{j}=\mu\end{subarray}}\frac{\nu!}{\prod_{j=1}^{\mu}m_{j}!}\,\prod_{j=1}^{\mu}Z_{j}(x,y)^{\,m_{j}}.$
Equivalently, the generating function identity holds:
$\left(\sum_{j\geq 1}Z_{j}(x,y)\,\tau^{j}\right)^{\nu}=\sum_{\mu\geq 0}\widehat{B}_{\mu,\nu}(Z_{1},\dots,Z_{\mu};x,y)\,\tau^{\mu},$
with the convention $\widehat{B}_{0,0}=1$ and $\widehat{B}_{\mu,\nu}=0$ when $\mu<\nu$. This definition follows the classical treatment of partial Bell polynomials (see Comtet (1974); Riordan (1958)), adapted here to the kernel notation.
Remark 2.5.
The constraint $\sum_{j=1}^{\mu}j\,m_{j}=\mu$ decomposes the integer $\mu$ into parts of sizes $j$ counted $m_{j}$ times, and $\sum_{j=1}^{\mu}m_{j}=\nu$ prescribes that exactly $\nu$ parts are used.
2.2. Non-Hilbert–Schmidt solution and its property
Theorem 2.6.
Let $K\in\mathbb{K}$. Then the eigenvalue equation (5) admits a positive spectral value $\lambda_{0}\in\mathbb{R}_{+}$ with a nontrivial eigenfunction $w(\cdot,y;\lambda_{0})\in L^{1}(\Omega^{d})$; moreover, $\lambda_{0}$ is simple. Fix $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$ and normalize $w$ so that $c_{0}(x_{0},y_{0})>0$. Under this normalization, $w$ admits the following uniformly convergent series representation:
(6) $w(x,y,x_{0},y_{0};\lambda_{0})=c_{0}(x_{0},y_{0})\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,n}}\,\Gamma_{n}(x,y,x_{0},y_{0}),$
where $\Gamma_{1}(x,y,x_{0},y_{0})=K^{(1)}(x,y)$ and, for $n\geq 2$,
(7) $\Gamma_{n}(x,y,x_{0},y_{0})=K^{(n)}(x,y)+\sum_{\ell=1}^{n-1}(-1)^{\ell}\sum_{k=\ell}^{n-1}K^{(n-k)}(x,y)\,\widehat{B}_{k,\ell}\!\Big(K^{(1)},K^{(2)},\dots,K^{(k)};x_{0},y_{0}\Big).$
We prove Theorem 2.6 by extending the following proposition, which asserts that (6) holds for kernels in $\mathbb{K}_{2}$.
Proposition 2.7.
If $K\in\mathbb{K}_{2}$, then the conclusion of Theorem 2.6 is valid.
The proof of Proposition 2.7 proceeds in several steps. In the first step, we represent the solution of (5) using the Fredholm determinant and the Neumann series for the resolvent of $\mathbf{K}\in\mathbb{K}_{2}$. Recall that the Fredholm determinant is well-defined for kernels in $\mathbb{K}_{2}$, though this is not necessarily the case for all kernels in $\mathbb{K}$. In the second step, we obtain the explicit representation (6)–(7) via Fredholm’s method, building on the first step. In the third step, we establish the correspondence between the Fredholm determinant expansion and the series in terms of iterated kernels. We also investigate how the operator norm of $\mathbf{K}$ determines the radius of convergence for both the Neumann series of the resolvent and the series in (6). Finally, we prove the existence of a dominant, simple spectral value $\lambda_{0}$. Notably, the arguments in the last two steps remain valid for general kernels in $\mathbb{K}$.
Step 1.
We begin the proof by analyzing the solution of (5) within the framework of Fredholm theory. First, we assume that $K\in\mathbb{K}_{2}$, for which the Fredholm determinant is well-defined and given by
(8) $D(\lambda)$ $=1+\sum_{m=1}^{\infty}\frac{1}{m!}\left(-\frac{1}{\lambda}\right)^{m}\int_{\Omega^{d}}\!\!\cdots\!\int_{\Omega^{d}}$
$\quad\times\det\!\begin{pmatrix}K(x_{1},x_{1})&K(x_{1},x_{2})&\cdots&K(x_{1},x_{m})\\ K(x_{2},x_{1})&K(x_{2},x_{2})&\cdots&K(x_{2},x_{m})\\ \vdots&\vdots&\ddots&\vdots\\ K(x_{m},x_{1})&K(x_{m},x_{2})&\cdots&K(x_{m},x_{m})\end{pmatrix}\prod_{k=1}^{m}dx_{k}.$
The resolvent of $\mathbf{K}$ admits the Neumann series representation:
(9) $\frac{1}{\lambda}\left(\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\right)^{-1}K\right)(x,y)=\sum_{n=1}^{\infty}\frac{1}{\lambda^{n}}K^{(n)}(x,y).$
Since $K(x,\cdot)\in L^{2}(\Omega^{d})$, there exists $M>0$ such that $K(x,y)\leq M$ for all $y\in\Omega^{d}$. Then $K^{(n)}(x,y)=\mathbf{K}^{\,n-1}K(x,y)\leq MN_{0}^{n-1}$, and hence
(10) $\frac{1}{\lambda}\left(\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\right)^{-1}K\right)(x,y)\leq\sum_{n=1}^{\infty}\frac{MN_{0}^{n-1}}{\lambda^{n}},$
which shows that the Neumann series (9) converges uniformly in $x\in\Omega^{d}$ for $|\lambda|>N_{0}$.
We define
(11) $D(x,y;\lambda):=D(\lambda)\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\right)^{-1}\frac{K(x,y)}{\lambda},$
which satisfies (5) at $\lambda=\lambda_{0}$ by Fredholm theory. Although $D(\lambda)\to 0$ and the resolvent diverges as $\lambda\downarrow\lambda_{0}$, their product remains finite, so (11) yields a valid solution at $\lambda_{0}$.
For $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$, introduce
(12) $w(x,y,x_{0},y_{0};\lambda):=c(x_{0},y_{0};\lambda)D(x,y;\lambda),$
with
(13) $c(x_{0},y_{0};\lambda):=\frac{c_{0}(x_{0},y_{0})}{D(\lambda)+D(x_{0},y_{0};\lambda)},$
where $c_{0}(x_{0},y_{0})$ is a fixed nonzero constant and we assume $\lim_{\lambda\downarrow\lambda_{0}}D(x_{0},y_{0};\lambda)\neq 0$. Combining (12) and (13) yields
(14) $\frac{c(x_{0},y_{0};\lambda)}{c_{0}(x_{0},y_{0})}D(\lambda)=1-\frac{w(x_{0},y_{0},x_{0},y_{0};\lambda)}{c_{0}(x_{0},y_{0})},$
which links the Fredholm determinant with the solution and serves as the basis for constructing the sequence $\Gamma_{n}$ in Step 2.
Step 2.
In this step, we establish the relation between the Fredholm determinant $D(\lambda)$ and the kernel evaluated at a fixed reference point $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$. Using (11) and (12), we first obtain
(15) $\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\right)w(x,y,x_{0},y_{0};\lambda)$ $=c(x_{0},y_{0};\lambda)D(\lambda)\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\right)\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\right)^{-1}\frac{K(x,y)}{\lambda}$
$=c(x_{0},y_{0};\lambda)D(\lambda)\frac{K(x,y)}{\lambda}.$
This leads to a reformulated version of the integral equation:
(16) $w(x,y,x_{0},y_{0};\lambda)$ $=c(x_{0},y_{0};\lambda)D(\lambda)\frac{K(x,y)}{\lambda}+\frac{1}{\lambda}\mathbf{K}w(x,y,x_{0},y_{0};\lambda)$
$=c(x_{0},y_{0};\lambda)D(\lambda)\frac{K(x,y)}{\lambda}+\int_{\Omega^{d}}\frac{K(x,\xi)}{\lambda}w(\xi,y,x_{0},y_{0};\lambda)\,d\xi.$
Following the original method of Fredholm (Fredholm, 1903), which is based on a formal series expansion of the kernel through successive iterations, we expand the solution $w$ into a series normalized by the nonzero constant $c_{0}(x_{0},y_{0})$:
(17) $w(x,y,x_{0},y_{0};\lambda)=\sum_{n=1}^{\infty}\frac{c_{0}(x_{0},y_{0})}{\lambda^{n}}\Gamma_{n}(x,y,x_{0},y_{0}).$
Assuming that the series (17) converges uniformly, the relation (14) together with (17) yields the following spectral identity:
(18) $1=\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\Gamma_{n}(x_{0},y_{0};x_{0},y_{0}),$
where $\Gamma_{1}(x_{0},y_{0};x_{0},y_{0})=K(x_{0},y_{0})$ by definition.
To explicitly determine the sequence $\Gamma_{n}$, substitute (14) and (17) into (16) to obtain
(19) $\sum_{n=1}^{\infty}\frac{1}{\lambda^{n}}\Gamma_{n}(x,y;x_{0},y_{0})$ $=\frac{1}{\lambda}K(x,y)+\sum_{n=1}^{\infty}\frac{1}{\lambda^{n+1}}\int_{\Omega^{d}}K(x,\xi)\Gamma_{n}(\xi,y;x_{0},y_{0})\,d\xi$
$\quad-\sum_{n=1}^{\infty}\frac{1}{\lambda^{n+1}}K(x,y)\Gamma_{n}(x_{0},y_{0};x_{0},y_{0}).$
Comparing coefficients of powers of $\lambda^{-1}$ on both sides of (19) gives the recurrence relation
(20) $\Gamma_{1}(x,y;x_{0},y_{0})$ $=K(x,y),$
$\Gamma_{n+1}(x,y;x_{0},y_{0})$ $=\int_{\Omega^{d}}K(x,\xi)\Gamma_{n}(\xi,y;x_{0},y_{0})\,d\xi-K(x,y)\Gamma_{n}(x_{0},y_{0};x_{0},y_{0}),\qquad n\geq 1.$
Step 3.
We first note that $c(x_{0},y_{0};\lambda)$ is a nonzero scalar for any fixed reference pair $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$. At this stage, we show that the explicit formula (7) indeed yields a solution to the Fredholm equation (17).
A reference-point operator and a Bell-polynomial expansion
We introduce a reference-point operator $\mathbf{P}$ associated with a fixed reference pair $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$. The purpose of $\mathbf{P}$ is to encode a “taboo at a reference point” directly at the kernel level. Although point evaluation is not a bounded linear functional on $L^{1}$, it is well-defined for kernels (as functions) and will be used here purely as an algebraic device to organize the taboo-type iterates.
Definition 2.8 (Reference-point operator).
For the iterated kernels $K^{(n)}$, define the reference-point moments
(21) $b_{n}:=K^{(n)}(x_{0},y_{0}),\qquad n\geq 1.$
We rewrite (16) in operator form as
(22) $\left(\mathbf{I}-\frac{1}{\lambda}(\mathbf{K}-\mathbf{P})\right)w(\cdot,\cdot,x_{0},y_{0};\lambda)=\frac{1}{\lambda}c_{0}(x_{0},y_{0})\,K\qquad\text{in }\mathcal{X}.$
Substituting (14) into the left-hand side of (16) and using the identity
$\Bigl(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\Bigr)w(\cdot,\cdot,x_{0},y_{0};\lambda)=\frac{1}{\lambda}c_{0}(x_{0},y_{0})\,K-\frac{1}{\lambda}\mathbf{P}\,w(\cdot,\cdot,x_{0},y_{0};\lambda),$
(with $\mathbf{P}$ defined in Definition 2.1), we obtain (22).
Formally solving (22), we obtain the resolvent representation
(23) $w(\cdot,\cdot,x_{0},y_{0};\lambda)=\frac{1}{\lambda}c_{0}(x_{0},y_{0})\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{A}\right)^{-1}K,$
that is,
$w(x,y,x_{0},y_{0};\lambda)=\frac{1}{\lambda}c_{0}(x_{0},y_{0})\left[\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{A})\right)^{-1}K\right](x,y).$
Expanding (23) into a Neumann series yields
$w(x,y,x_{0},y_{0};\lambda)=c_{0}(x_{0},y_{0})\sum_{n=1}^{\infty}\frac{1}{\lambda^{n}}\,\bigl(\mathbf{A}^{n-1}K\bigr)(x,y).$
Definition 2.9 (Taboo-type iterates).
Define $\Gamma_{1}:=K^{(1)}=K$ and, for $n\geq 1$,
(24) $\Gamma_{n+1}:=\mathbf{A}\Gamma_{n}.$
Equivalently,
(25) $\Gamma_{n}=\mathbf{A}^{\,n-1}K,\qquad n\geq 1.$
Finally set
(26) $a_{n}:=\Gamma_{n}(x_{0},y_{0};x_{0},y_{0}),\qquad n\geq 1.$
Remark 2.10 (Why “taboo-type”).
At this stage, the terminology “taboo” is only suggestive: the iterates $\{\Gamma_{n}\}_{n\geq 1}$ are defined algebraically by the recursion $\Gamma_{n+1}=\mathbf{A}\Gamma_{n}$ and need not yet be interpreted as probabilities. The reason for the name is that, in the discrete-time (matrix/Markov-chain) setting studied in Section 3 (and in particular in the taboo probability formulation recalled there), the analogous “taboo” quantities are generated by a recursion of the same structural form: one starts from the original transition object and iterates a taboo-modified transition operator obtained by subtracting a rank-one term that reinjects mass through a distinguished state. Here, $\mathbf{P}$ plays exactly this algebraic role at the kernel level: it is a rank-one operator built from the reference pair $(x_{0},y_{0})$, and the operator $\mathbf{A}$ is the reference-point analogue of a taboo-modified transition operator. Section 3 makes this analogy explicit by identifying $\Gamma_{n}$ (and the moments $a_{n}:=\Gamma_{n}(x_{0},y_{0};x_{0},y_{0})$) with the continuous-state counterparts of taboo recursions and their associated renewal/characteristic relations.
We recall the ordinary partial Bell polynomials $\widehat{B}_{m,\ell}$ by the generating function identity
(27) $\left(\sum_{j\geq 1}Z_{j}\,\tau^{j}\right)^{\ell}=\sum_{m\geq 0}\widehat{B}_{m,\ell}(Z_{1},Z_{2},\dots)\,\tau^{m},\qquad\ell\geq 0,$
with the convention $\widehat{B}_{0,0}=1$ and $\widehat{B}_{m,\ell}=0$ for $m<\ell$.
Lemma 2.11 (Bell-polynomial representation of $\Gamma_{n}$).
Assume Definition 2.1 and Definition 2.1. Then, for every $n\geq 1$,
(28) $\Gamma_{n}=\sum_{k=1}^{n}K^{(k)}\sum_{\ell=0}^{n-k}(-1)^{\ell}\,\widehat{B}_{\,n-k,\ \ell}(b_{1},b_{2},\dots),$
and consequently the reference-point moments $a_{n}:=\Gamma_{n}(x_{0},y_{0},x_{0},y_{0})$ satisfy
(29) $a_{n}=\sum_{k=1}^{n}b_{k}\sum_{\ell=0}^{n-k}(-1)^{\ell}\,\widehat{B}_{\,n-k,\ \ell}(b_{1},b_{2},\dots,b_{n-k};x_{0},y_{0}).$
Proof.
The identities are purely algebraic: we expand $\Gamma_{n}=(\mathbf{K}-\mathbf{P})^{n-1}K$ and identify the resulting combinatorics with ordinary partial Bell polynomials.
Step 1: Word expansion of $(\mathbf{K}-\mathbf{P})^{n-1}K$. Iterating (25) gives, for $n\geq 1$,
$\Gamma_{n}=(\mathbf{K}-\mathbf{P})^{\,n-1}K.$
Expanding $\mathbf{A}^{\,n-1}$ yields a sum over words containing $\ell$ occurrences of $\mathbf{P}$ and n1n-1-$\ell$ occurrences of $\mathbf{K}$:
$\mathbf{A}^{\,n-1}=\sum_{\ell=0}^{n-1}(-1)^{\ell}\sum_{\begin{subarray}{c}i_{0},\dots,i_{\ell}\geq 0\\ i_{0}+\cdots+i_{\ell}=n-1-\ell\end{subarray}}\mathbf{K}^{i_{0}}\mathbf{P}\,\mathbf{K}^{i_{1}}\mathbf{P}\cdots\mathbf{P}\,\mathbf{K}^{i_{\ell}}.$
Applying this to $K$ gives
(30) $\Gamma_{n}=\sum_{\ell=0}^{n-1}(-1)^{\ell}\sum_{\begin{subarray}{c}i_{0},\dots,i_{\ell}\geq 0\\ i_{0}+\cdots+i_{\ell}=n-1-\ell\end{subarray}}\mathbf{K}^{i_{0}}\mathbf{P}\,\mathbf{K}^{i_{1}}\mathbf{P}\cdots\mathbf{P}\,\mathbf{K}^{i_{\ell}}K.$
Step 2: Collapse each word using the definition of $\mathbf{P}$. Since $K^{(1)}=K$ and $K^{(n+1)}=\mathbf{K}K^{(n)}$, we have $\mathbf{K}^{j}K=K^{(j+1)}$ for $j\geq 0$. Hence, for $j\geq 0$,
$\mathbf{P}\,\mathbf{K}^{j}K=K\,(\mathbf{K}^{j}K)(x_{0},y_{0})=K\,K^{(j+1)}(x_{0},y_{0})=K\,b_{j+1}.$
Evaluating the word in (30) from right to left yields
$\mathbf{K}^{i_{0}}\mathbf{P}\,\mathbf{K}^{i_{1}}\mathbf{P}\cdots\mathbf{P}\,\mathbf{K}^{i_{\ell}}K=K^{(i_{0}+1)}\prod_{r=1}^{\ell}b_{i_{r}+1}.$
Substituting into (30) we obtain
(31) $\Gamma_{n}=\sum_{\ell=0}^{n-1}(-1)^{\ell}\sum_{\begin{subarray}{c}i_{0},\dots,i_{\ell}\geq 0\\ i_{0}+\cdots+i_{\ell}=n-1-\ell\end{subarray}}K^{(i_{0}+1)}\prod_{r=1}^{\ell}b_{i_{r}+1}.$
Step 3: Identify the inner composition sum with $\widehat{B}_{m,\ell}$. Fix $n\geq 1$ and set $k:=i_{0}+1\in\{1,\dots,n\}$. Then m:=nkm:=n-k and the constraint in (31) becomes
$i_{1}+\cdots+i_{\ell}=(n-k)-\ell=m-\ell.$
Let $j_{r}:=i_{r}+1\geq 1$ for $r=1,\dots,\ell$. Then $j_{1}+\cdots+j_{\ell}=m$ and
$\prod_{r=1}^{\ell}b_{i_{r}+1}=\prod_{r=1}^{\ell}b_{j_{r}}.$
Therefore,
$\sum_{\begin{subarray}{c}i_{1},\dots,i_{\ell}\geq 0\\ i_{1}+\cdots+i_{\ell}=m-\ell\end{subarray}}\prod_{r=1}^{\ell}b_{i_{r}+1}=\sum_{\begin{subarray}{c}j_{1},\dots,j_{\ell}\geq 1\\ j_{1}+\cdots+j_{\ell}=m\end{subarray}}\prod_{r=1}^{\ell}b_{j_{r}}=\widehat{B}_{m,\ell}(b_{1},b_{2},\dots)$
by (27). Plugging this into (31) and recalling m=nkm=n-k yields (28). Evaluating at $(x_{0},y_{0})$ and using (21)–(26) gives (29). ∎
Remark 2.12.
The first few terms illustrate the structure:
$\Gamma_{1}=K^{(1)},\qquad\Gamma_{2}=K^{(2)}-b_{1}K^{(1)},\qquad\Gamma_{3}=K^{(3)}-b_{1}K^{(2)}+(b_{1}^{2}-b_{2})K^{(1)}.$
Corollary 2.13 (Existence of a focal state controlled by the operator norm).
Assume that $\Omega^{d}\subset\mathbb{R}^{d}$ is unbounded (hence $|\Omega^{d}|=\infty$), and let $K\in\mathbb{K}$ be strongly positive, i.e. $K(x,y)>0$ for all $(x,y)\in\Omega^{d}\times\Omega^{d}$. Let $\mathbf{K}:L^{1}(\Omega^{d})\to L^{1}(\Omega^{d})$ be the integral operator (2). Then there exists a focal state $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$ such that
$\|\mathbf{K}\|_{\mathrm{op}}\ \geq\ K(x_{0},y_{0}).$
Proof.
Since $K\geq 0$ and $K\in\mathbb{K}$, we have
$\|\mathbf{K}\|_{\mathrm{op}}=\operatorname*{ess\,sup}_{y\in\Omega^{d}}\int_{\Omega^{d}}K(x,y)\,dx<\infty.$
Suppose, to the contrary, that $K(x,y)>\|\mathbf{K}\|_{\mathrm{op}}$ for all $(x,y)\in\Omega^{d}\times\Omega^{d}$. Fix any $y\in\Omega^{d}$. Then
$\int_{\Omega^{d}}K(x,y)\,dx\ \geq\ \int_{\Omega^{d}}\|\mathbf{K}\|_{\mathrm{op}}\,dx=\|\mathbf{K}\|_{\mathrm{op}}\ |\Omega^{d}|\ =\ \infty,$
because $\Omega^{d}$ is unbounded and thus $|\Omega^{d}|=\infty$. This contradicts the finiteness of $\operatorname*{ess\,sup}_{y}\int_{\Omega^{d}}K(x,y)\,dx$. Hence there exists $(x_{0},y_{0})$ with $K(x_{0},y_{0})\leq\|\mathbf{K}\|_{\mathrm{op}}$, which is the desired inequality. ∎
Lemma 2.14 (Spectral structure of the reference-point operator).
Fix $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$ that gives $K(x_{0},y_{0})\leq N_{0}$ and define $\mathbf{P}$ by
$(\mathbf{P}F)(x,y):=K(x,y)\,F(x_{0},y_{0}).$
Set $K_{0}:=K(x_{0},y_{0})>0$. Then:
  1. (1)
    $\mathrm{Ran}(\mathbf{P})=\mathrm{span}\{K\}$ and $\mathbf{P}$ has rank one.
  2. (2)
    $\mathbf{P}K=K_{0}\,K$, hence $K$ is an eigenfunction with eigenvalue $K_{0}$.
  3. (3)
    For every $n\geq 1$, $\mathbf{P}^{\,n}=K_{0}^{\,n-1}\mathbf{P}$. Consequently, the spectrum of $\mathbf{P}$ is
    $\sigma(\mathbf{P})=\{0,\,K_{0}\},$
    and the spectral radius satisfies $\rho(\mathbf{P})=K_{0}$. In particular, the (unique) positive eigenvalue $K_{0}$ is the maximal eigenvalue.
Remark 2.15 (Notation: range and span).
For a linear operator $\mathbf{T}$ on a vector space $\mathcal{V}$, we write
$\mathrm{Ran}(\mathbf{T}):=\{\mathbf{T}v:\ v\in\mathcal{V}\}$
for the range (image) of $\mathbf{T}$. For $v\in\mathcal{V}$, we write
$\mathrm{span}\{v\}:=\{\alpha v:\ \alpha\in\mathbb{R}\}$
for the one-dimensional linear subspace generated by $v$ (and similarly $\mathrm{span}\{v_{1},\dots,v_{m}\}:=\{\sum_{i=1}^{m}\alpha_{i}v_{i}:\alpha_{i}\in\mathbb{R}\}$).
Proof.
(1) By definition, $\mathbf{P}F$ is always a scalar multiple of $K$, hence $\mathrm{Ran}(\mathbf{P})\subset\mathrm{span}\{K\}$. Since $\mathbf{P}(1)=K$, equality holds and $\mathbf{P}$ has rank one.
(2) Taking $F=K$ gives
$(\mathbf{P}K)(x,y)=K(x,y)\,K(x_{0},y_{0})=K_{0}\,K(x,y).$
(3) Using $\mathbf{P}F=K\,F(x_{0},y_{0})$ and evaluating at $(x_{0},y_{0})$,
$(\mathbf{P}F)(x_{0},y_{0})=K_{0}\,F(x_{0},y_{0}).$
Hence
$\mathbf{P}^{2}F=\mathbf{P}\big(K\,F(x_{0},y_{0})\big)=K\,(\mathbf{P}F)(x_{0},y_{0})=K\,K_{0}\,F(x_{0},y_{0})=K_{0}\,\mathbf{P}F,$
and inductively $\mathbf{P}^{\,n}=K_{0}^{\,n-1}\mathbf{P}$. Thus the only possible nonzero spectral value is $K_{0}$, which is indeed an eigenvalue by (2). Since $\mathbf{P}$ has rank one, all remaining spectral values are 0. Therefore $\sigma(\mathbf{P})=\{0,K_{0}\}$ and $\rho(\mathbf{P})=K_{0}$. ∎
We now aim to control the convergence of the series (17) by selecting a reference pair $(x_{0},y_{0})$ for which convergence is guaranteed. To rigorously justify (17), however, it is necessary to verify that its radius of convergence is strictly greater than that of the Neumann series in (9).
Lemma 2.16.
Let $K\in\mathbb{K}$ be a positive continuous kernel on $\Omega^{d}\times\Omega^{d}$. Assume that the Neumann series for $\mathbf{K}$ has radius of convergence $R>0$. Then there exists a reference pair $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$ such that, with $K_{0}:=K(x_{0},y_{0})$,
$0\leq N_{0}-K_{0}<\frac{1}{R}.$
Moreover, one may choose $(x_{0},y_{0})$ so that $K_{0}<N_{0}$ whenever $K$ is not constant. (If $K$ is constant, then $N_{0}-K_{0}=0$ for every pair.)
Proof.
By the definition of the supremum, for $\varepsilon:=1/R$ there exists $(x_{0},y_{0})$ such that $K(x_{0},y_{0})>N_{0}-\varepsilon$, hence $0\leq N_{0}-K_{0}<1/R$. If $K$ is not constant, there also exist points where $K<N_{0}$, so we may ensure $K_{0}<N_{0}$ while keeping $K_{0}$ arbitrarily close to $N_{0}$. ∎
Key point. By choosing the reference pair $(x_{0},y_{0})$ so that $N_{0}-K(x_{0},y_{0})$ is small, the rank-one subtraction $\mathbf{K}-\mathbf{P}$ becomes a strictly “smaller” transition operator, and the resulting genealogical series can be made to converge at the dominant spectral value.
By Lemma 2.16, we may choose $(x_{0},y_{0})$ so that
$N_{0}-\frac{1}{R}<K_{0}\leq N_{0}.$
In particular, unless $K$ is constant, we can ensure $K_{0}<N_{0}$ while making $N_{0}-K_{0}$ arbitrarily small. Let $M_{0}$ be $K(x,y)\leq M_{0}$ The radius of convergence for the series (17) can be determined as follows:
(32) $\begin{bmatrix}\Gamma_{n}(x,y;x_{0},y_{0})\\[2.84526pt] a_{n}\end{bmatrix}$ $\leq\begin{bmatrix}N_{0}&-K_{0}\\ 0&N_{0}-K_{0}\end{bmatrix}^{n-1}\begin{bmatrix}M_{0}\\ M_{0}\end{bmatrix}=\begin{bmatrix}M_{0}(N_{0}-K_{0})^{n-1}\\ M_{0}(N_{0}-K_{0})^{n-1}\end{bmatrix},\qquad n\geq 1.$
Indeed, $\Gamma_{1}=K\leq M_{0}$ and $a_{1}=K_{0}\leq N_{0}$, so the initial vector can be bounded by $(M_{0},N_{0})^{\top}$.
Consequently, the series (17) converges whenever
$|\lambda|>N_{0}-K_{0}.$
By choosing $(x_{0},y_{0})$ so that $N_{0}-K_{0}<1/R$, the convergence region strictly contains that of the Neumann series in (9), and in particular includes the dominant spectral value $\lambda_{0}$.
Adjoint counterpart. In applications (Sections 34), the adjoint eigenfunction plays the role of a reproductive value (dual eigenfunction). For completeness, we record the adjoint analogue of the reference-point expansion, which has the same recursion structure as the primal construction.
Remark 2.17 (Adjoint expansion).
We briefly outline the analogous expansion for the adjoint operator. Let $\mathbf{K}^{*}$ act on functions $v(y,\cdot)\in L^{\infty}(\Omega^{d})$ by
$(\mathbf{K}^{*}v)(y,x):=\int_{\Omega^{d}}v(y,\xi)\,K(\xi,x)\,d\xi.$
Under the assumptions of Theorem 2.6, consider the adjoint eigen-equation
(33) $v(y,x;\lambda)=\frac{1}{\lambda}\,(\mathbf{K}^{*}v)(y,x;\lambda)=\frac{1}{\lambda}\int_{\Omega^{d}}v(y,\xi;\lambda)\,K(\xi,x)\,d\xi,\qquad v(y,\cdot;\lambda)\in L^{\infty}(\Omega^{d}).$
Fix the same reference pair $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$ and introduce the transpose kernel
$K^{\top}(x,y):=K(y,x)\qquad\text{(so that $K^{\top}(\cdot,\cdot)$ is a kernel on $\Omega^{d}\times\Omega^{d}$).}$
Define the adjoint reference-point operator $\mathbf{P}^{*}$ by
(34) $(\mathbf{P}^{*}G)(y,x):=G(y_{0},x_{0})\,K^{\top},\qquad G\in\mathcal{X}^{*}.$
Then $\mathbf{P}^{*}$ is rank one and plays the same algebraic role as $\mathbf{P}$.
Define the adjoint taboo-type iterates $\{\Gamma_{n}^{*}\}_{n\geq 1}$ by
(35) $\Gamma_{1}^{*}:=K^{\top},\qquad\Gamma_{n+1}^{*}:=(\mathbf{K}^{*}-\mathbf{P}^{*})\Gamma_{n}^{*},\qquad n\geq 1,$
equivalently,
$\Gamma_{n}^{*}=(\mathbf{K}^{*}-\mathbf{P}^{*})^{\,n-1}K^{\top},\qquad n\geq 1.$
In kernel form, (35) reads
(36) $\Gamma_{n+1}^{*}(y_{0},x_{0};y,x)=\int_{\Omega^{d}}\Gamma_{n}^{*}(y_{0},x_{0};y,\xi)\,K(\xi,x)\,d\xi-\Gamma_{n}^{*}(y_{0},x_{0};y_{0},x_{0})\,K^{\top}(x,y),\qquad n\geq 1.$
Consequently, the solution admits the resolvent/Neumann expansion
(37) $v(y,x;\lambda)=\frac{c_{1}(x_{0},y_{0})}{\lambda}\left[\left(\mathbf{I}-\frac{1}{\lambda}(\mathbf{K}^{*}-\mathbf{P}^{*})\right)^{-1}K^{\top}\right](x,y)=c_{1}(x_{0},y_{0})\sum_{n=1}^{\infty}\frac{1}{\lambda^{n}}\,\Gamma_{n}^{*}(y_{0},x_{0};y,x),$
with $c_{1}(x_{0},y_{0})\neq 0$. In particular, evaluating at $\lambda=\lambda_{0}$ yields an explicit representation of the adjoint eigenfunction.
In the next step, we complete the proof of Proposition 2.7 by characterizing the spectral value $\lambda_{0}$.
Step 4.
We now establish the existence of a spectral value $\lambda_{0}$ as assumed in Theorem 2.6. As observed in (14), any $\lambda_{0}$ satisfying (18) corresponds to a zero of the Fredholm determinant, i.e., $D(\lambda_{0})=0$. Furthermore, we show that $\lambda_{0}$ is both dominant and simple, as stated in the following proposition, which completes the proof of Proposition 2.7.
Proposition 2.18.
There exists a dominant spectral value $\lambda_{0}$ satisfying (18) that is positive, real, simple, and coincides with the spectral radius.
Alternative proof of Proposition 2.18.
Fix a reference pair $(x_{0},y_{0})\in\Omega^{d}\times\Omega^{d}$ as in Lemma 2.16 and set
$K_{0}:=K(x_{0},y_{0})>0.$
Recall $a_{n}:=\Gamma_{n}(x_{0},y_{0};x_{0},y_{0})$ and define, for $\lambda>|N_{0}-K_{0}|$,
$F(\lambda):=\sum_{n=1}^{\infty}\frac{a_{n}}{\lambda^{n}},\qquad\mathcal{D}(\lambda):=1-F(\lambda).$
Step 1 (Generating-function identity). Let $\rho(\mathbf{K})$ be the spectral radius of $\mathbf{K}.$ Then for every $\lambda>\rho(\mathbf{K})$, the resolvent exists and we have
(38) $F(\lambda)=\frac{\bigl((\lambda\mathbf{I}-\mathbf{K})^{-1}K\bigr)(x_{0},y_{0})}{1+\bigl((\lambda\mathbf{I}-\mathbf{K})^{-1}K\bigr)(x_{0},y_{0})}.$
Indeed, for $\lambda>\rho(\mathbf{K})$ the Neumann expansion $(\lambda\mathbf{I}-\mathbf{K})^{-1}=\lambda^{-1}\sum_{m\geq 0}(\mathbf{K}/\lambda)^{m}$ holds, and combining it with (16) (equivalently (22)–(23)) yields (38).
Step 2: Existence of $\lambda_{0}\in(|N_{0}-K_{0}|,\infty)$ with $F(\lambda_{0})=1$.
By the assumed resolvent bound,
$F(\lambda)=\Bigl[\Bigl(\mathbf{I}-\frac{1}{\lambda}(\mathbf{K}-\mathbf{P})\Bigr)^{-1}K\Bigr](x_{0},y_{0})\leq\frac{N_{0}}{\lambda-|N_{0}-K_{0}|},\qquad\lambda>|N_{0}-K_{0}|.$
Hence $F(\lambda)<1$ for every $\lambda>\lambda_{+}:=|N_{0}-K_{0}|+N_{0}$, and therefore $\mathcal{D}(\lambda_{+})>0$.
On the other hand, since $a_{1}=\Gamma_{1}(x_{0},y_{0};x_{0},y_{0})=K_{0}$, we have
$F(\lambda)\geq\frac{a_{1}}{\lambda}=\frac{K_{0}}{\lambda},\qquad\lambda>|N_{0}-K_{0}|.$
By Lemma 2.16, we may choose $(x_{0},y_{0})$ so that $|N_{0}-K_{0}|<K_{0}$. Fix any $\lambda_{-}\in(|N_{0}-K_{0}|,K_{0})$. Then $F(\lambda_{-})\geq K_{0}/\lambda_{-}>1$, i.e., $\mathcal{D}(\lambda_{-})<0$. Since $|a_{n}|$ admits the exponential bound obtained in (32), the series defining $F(\lambda)$ converges absolutely for $\lambda>|N_{0}-K_{0}|$, and hence $F$ is continuous on $(|N_{0}-K_{0}|,\infty)$. Hence, by the intermediate value theorem, there exists $\lambda_{0}\in(\lambda_{-},\lambda_{+})$ such that $\mathcal{D}(\lambda_{0})=0$, equivalently $F(\lambda_{0})=1$. This is precisely (18).
Step 3: Monotonicity on the resolvent region. Let
$S(\lambda):=\bigl((\lambda\mathbf{I}-\mathbf{K})^{-1}K\bigr)(x_{0},y_{0}),\qquad\lambda>\rho(\mathbf{K}).$
Since $\mathbf{K}$ is positive, $(\lambda\mathbf{I}-\mathbf{K})^{-1}$ is also positive for $\lambda>\rho(\mathbf{K})$, hence $S(\lambda)>0$ there. Moreover, for $\lambda_{2}>\lambda_{1}>\rho(\mathbf{K})$, the resolvent identity gives
$(\lambda_{1}\mathbf{I}-\mathbf{K})^{-1}-(\lambda_{2}\mathbf{I}-\mathbf{K})^{-1}=(\lambda_{2}-\lambda_{1})(\lambda_{1}\mathbf{I}-\mathbf{K})^{-1}(\lambda_{2}\mathbf{I}-\mathbf{K})^{-1}\geq 0,$
so $(\lambda\mathbf{I}-\mathbf{K})^{-1}$ is decreasing in $\lambda$ in the operator order, and therefore $S(\lambda)$ is decreasing in $\lambda$. Since the identity (38) implies that $F(\lambda)$ is strictly decreasing on $(\rho(\mathbf{K}),\infty)$ (hence $\mathcal{D}(\lambda)$ is strictly increasing there). In particular, there can be no more than one solution of $F(\lambda)=1$ in the resolvent region $(\rho(\mathbf{K}),\infty)$. Since the scalar map $u\mapsto u/(1+u)$ is strictly increasing on $(0,\infty)$, (38) implies that $F(\lambda)$ is strictly decreasing on $(\rho(\mathbf{K}),\infty)$. In particular, the equation $F(\lambda)=1$ has at most one solution in the resolvent region $(\rho(\mathbf{K}),\infty)$.
Furthermore, strict monotonicity implies that $\mathcal{D}(\lambda)=1-F(\lambda)$ is strictly increasing near $\lambda_{0}$, hence the root is simple.
Step 4: $\lambda_{0}$ is a spectral value of $\mathbf{K}$. Rearranging the Fredholm equation (16) gives
$\Bigl(\mathbf{I}-\frac{1}{\lambda}\mathbf{K}\Bigr)w(\cdot,\cdot,x_{0},y_{0};\lambda)=c(x_{0},y_{0};\lambda)\,\mathcal{D}(\lambda)\,\frac{K}{\lambda}\qquad\text{in }\mathcal{X}.$
By construction, for $\lambda>|N_{0}-K_{0}|$,
$w(x,y,x_{0},y_{0};\lambda)=c_{0}(x_{0},y_{0})\sum_{n=1}^{\infty}\frac{1}{\lambda^{n}}\,\Gamma_{n}(x,y;x_{0},y_{0}),$
so at $\lambda=\lambda_{0}$ the series converges and $w(\cdot;\lambda_{0})\not\equiv 0$. Since $\mathcal{D}(\lambda_{0})=0$, substituting $\lambda=\lambda_{0}$ yields
$\Bigl(\mathbf{I}-\frac{1}{\lambda_{0}}\mathbf{K}\Bigr)w(\cdot;\lambda_{0})=0,\quad\text{i.e.,}\quad\mathbf{K}\,w(\cdot;\lambda_{0})=\lambda_{0}\,w(\cdot;\lambda_{0}).$
Thus $\lambda_{0}\in\sigma(\mathbf{K})$.
Step 5: Dominance and $\lambda_{0}=\rho(\mathbf{K})$. Since $\mathbf{K}$ is strongly positive, $\lambda_{0}>0$ admits a strictly positive eigenfunction, and therefore $\lambda_{0}$ is the dominant spectral value. In particular, $\lambda_{0}=\rho(\mathbf{K})$, and it is simple. ∎
2.3. Proof of Theorem 2.6
Proof.
From Lemma 2.16, the series (6) converges uniformly. Since the integral kernels appearing in the expansion (17) and in equation (18) belong to the class $\mathbb{K}$, they satisfy the required regularity and decay conditions. Moreover, Proposition 2.18 ensures the existence of the spectral value $\lambda_{0}$ for kernels in $\mathbb{K}$. Combining these results completes the proof of Theorem 2.6. ∎
In this paper, we refer to the solution (6) of the Fredholm integral equation with kernel $K\in\mathbb{K}$ on the spectral set as a non–Hilbert–Schmidt solution.
3. Application to simple integral projection models at discrete time
In this section, we investigate the eigenvalue problem for the simplest IPM, using the eigenfunction representation provided by Theorem 2.6. This representation also offers a Markovian and biological interpretation of the system.
Let $P_{t}(x)$ denote the cohort at state $x$ at time $t$. The simplest IPM considered here is given by
(39) $P_{t+1}\left(x\right)=\int_{\Omega^{d}}K\left(x,y\right)P_{t}\left(y\right)\,dy,\quad P_{0}\left(x\right)\in L^{1}\left(\Omega^{d}\right),\quad P_{0}\left(x\right)\geq 0,$
where the integral kernel $K\left(\cdot\right)$ is defined in Theorem 2.6.
Before constructing the eigenvalue problem for this IPM, we introduce one convention. The representation used in Theorem 2.6 involves three points not directly related to the integral operator, namely $y$, $x_{0}$, and $y_{0}$. These points can be chosen freely provided that the chosen reference pair satisfies the condition of Lemma 2.16. However, in consideration of the usefulness of biological interpretation, which will be discussed later, we therefore impose the convention $y=x_{0}=y_{0}$ and choose the reference point $y_{0}$ so that the associated pair $(y_{0},y_{0})$ satisfies Lemma 2.16. Based on these assumptions, we will omit variables from this chapter as follows:
$w_{0}\left(x,y_{0}\right)$ $=w_{0}\left(x,y_{0},y_{0},y_{0}\right),$
$v_{0}\left(y_{0},x\right)$ $=v_{0}\left(y_{0},y_{0},y_{0},x\right),$
$\Gamma_{n}(x,y_{0})$ $=\Gamma_{n}(x,y_{0};y_{0},y_{0}),$
$\Gamma_{n}^{*}(y_{0},x)$ $=\Gamma_{n}^{*}(y_{0},y_{0};y_{0},x).$
Under this convention, the eigenfunction $w_{0}(x,y_{0})$ corresponding to (39) is expressed as the following series:
(40) $w_{0}\left(x,y_{0}\right)$ $=c_{w}\left(y_{0}\right)\left(\sum_{n=1}^{\infty}\frac{\Gamma_{n}\left(x,y_{0}\right)}{\lambda_{0}^{n}}\right),\quad c_{w}\left(y_{0}\right)\neq 0,$
(41) $\Gamma_{1}\left(x,y\right)$ $=K\left(x,y_{0}\right),$
(42) $\Gamma_{n+1}\left(x,y_{0}\right)$ $=\int_{\Omega^{d}}K\left(x,\xi\right)\Gamma_{n}\left(\xi,y\right)-K\left(x,y_{0}\right)\Gamma_{n}\left(y_{0},y_{0}\right)\,d\xi,\qquad n\geq 1.$
For the adjoint eigenfunction $v_{0}(y,x)$, we adopt the same values for the three points $(y=x_{0}=y_{0})$ in (40), yielding
(43) $v_{0}\left(y_{0},x\right)$ $=c_{v}\left(y\right)\left(\sum_{n=1}^{\infty}\frac{\Gamma_{n}^{*}\left(y_{0},x\right)}{\lambda_{0}^{n}}\right),\quad c_{v}\left(y_{0}\right)\neq 0,$
(44) $\Gamma_{1}^{*}\left(y_{0},x\right)$ $=K\left(y_{0},x\right),$
(45) $\Gamma_{n+1}^{*}\left(y_{0},x\right)$ $=\int_{\Omega^{d}}\Gamma_{n}^{*}\left(y_{0},\eta\right)K\left(\eta,x\right)-\Gamma_{n}^{*}\left(y_{0},y_{0}\right)K\left(y_{0},x\right)\ d\eta,\qquad n\geq 1.$
Note that, for notational consistency, the variables $x$ and $y$ in (43) have been interchanged to align with the convention of expressing functions with respect to $x$.
3.1. Asymptotic characterization by the eigensystem
Recall that the cohort dynamics (39) can be written as
$P_{t}=\mathbf{K}^{\,t}P_{0}\qquad(t\in\mathbb{N}),$
where $\mathbf{K}:L^{1}(\Omega^{d})\to L^{1}(\Omega^{d})$ is the integral operator
$(\mathbf{K}f)(x):=\int_{\Omega^{d}}K(x,y)f(y)\,dy.$
The dual pairing is
(46) $\langle f,g\rangle_{x}:=\int_{\Omega^{d}}f(x)g(x)\,dx,\qquad f\in L^{\infty}(\Omega^{d}),\ g\in L^{1}(\Omega^{d}),$
and the adjoint operator $\mathbf{K}^{*}:L^{\infty}(\Omega^{d})\to L^{\infty}(\Omega^{d})$ is given by
$(\mathbf{K}^{*}v)(y)=\int_{\Omega^{d}}v(x)\,K(x,y)\,dx.$
Theorem 3.1 (Asymptotics of the cohort).
Assume that Proposition 2.18 holds and that the dominant spectral value $\lambda_{0}=\rho(\mathbf{K})$ is isolated in the spectrum of $\mathbf{K}$ in the sense that there exists $\theta\in(0,1)$ such that
$\sigma(\mathbf{K})\setminus\{\lambda_{0}\}\subset\{z\in\mathbb{C}:\ |z|\leq\theta\lambda_{0}\}.$
Let $w_{0}(\cdot,y)\in L^{1}(\Omega^{d})$ and $v_{0}(y,\cdot)\in L^{\infty}(\Omega^{d})$ be nontrivial eigenfunctions satisfying
$\mathbf{K}\,w_{0}(\cdot,y_{0})=\lambda_{0}\,w_{0}(\cdot,y_{0}),\qquad\mathbf{K}^{*}v_{0}(y_{0},\cdot)=\lambda_{0}\,v_{0}(y_{0},\cdot),$
with $\langle v_{0},w_{0}\rangle_{x}\neq 0$. Then there exist constants $C>0$ and $\delta>0$ such that, for every $P_{0}\in L^{1}(\Omega^{d})$ with $P_{0}\geq 0$,
(47) $P_{t}(x)=\frac{\langle v_{0},P_{0}\rangle_{x}}{\langle v_{0},w_{0}\rangle_{x}}\,\lambda_{0}^{t}\,w_{0}(x,y_{0})\Bigl(1+O(e^{-\delta t})\Bigr),\qquad t\to\infty,$
where $\langle\cdot,\cdot\rangle_{x}$ is defined in (46).
Proof.
Let $\Gamma$ be a positively oriented circle in $\mathbb{C}$ centered at $\lambda_{0}$ that encloses no other point of $\sigma(\mathbf{K})$. Define the Riesz projection
$\mathbf{\Pi}:=\frac{1}{2\pi i}\oint_{\Gamma}(z\mathbf{I}-\mathbf{K})^{-1}\,dz\quad\text{on }L^{1}(\Omega^{d}).$
Then $\mathbf{\Pi}$ is a bounded projection commuting with $\mathbf{K}$ and satisfying $\mathbf{K}\mathbf{\Pi}=\lambda_{0}\mathbf{\Pi}$.
Since $\lambda_{0}$ is simple, $\mathrm{Ran}(\mathbf{\Pi})=\mathrm{span}\{w_{0}(\cdot,y)\}$. Hence there exists a bounded linear functional $\ell$ on $L^{1}(\Omega^{d})$ such that
$\mathbf{\Pi}f=\ell(f)\,w_{0}(\cdot,y_{0}),\qquad f\in L^{1}(\Omega^{d}).$
Using the adjoint eigenfunction, we compute $\ell$ as follows. Because $\mathbf{\Pi}$ commutes with $\mathbf{K}$, its adjoint $\mathbf{\Pi}^{*}$ commutes with $\mathbf{K}^{*}$. Moreover, $\mathbf{K}^{*}v_{0}=\lambda_{0}v_{0}$ implies $\mathbf{\Pi}^{*}v_{0}=v_{0}$. Therefore, for every $f\in L^{1}(\Omega^{d})$,
$\langle v_{0},\mathbf{\Pi}f\rangle_{x}=\langle\mathbf{\Pi}^{*}v_{0},f\rangle_{x}=\langle v_{0},f\rangle_{x}.$
On the other hand, $\mathbf{\Pi}f=\ell(f)w_{0}$ gives
$\langle v_{0},\mathbf{\Pi}f\rangle_{x}=\ell(f)\,\langle v_{0},w_{0}\rangle_{x}.$
Hence
$\ell(f)=\frac{\langle v_{0},f\rangle_{x}}{\langle v_{0},w_{0}\rangle_{x}},\qquad\mathbf{\Pi}f=\frac{\langle v_{0},f\rangle_{x}}{\langle v_{0},w_{0}\rangle_{x}}\,w_{0}.$
Now decompose $\mathbf{K}$ as
$\mathbf{K}=\lambda_{0}\mathbf{\Pi}+\mathbf{R},\qquad\mathbf{R}:=\mathbf{K}(\mathbf{I}-\mathbf{\Pi}).$
Then $\mathbf{\Pi}\mathbf{R}=\mathbf{R}\mathbf{\Pi}=0$ and
$\mathbf{K}^{t}=\lambda_{0}^{t}\mathbf{\Pi}+\mathbf{R}^{t}\qquad(t\in\mathbb{N}).$
By the spectral gap assumption, $\sigma(\mathbf{R})\subset\sigma(\mathbf{K})\setminus\{\lambda_{0}\}$ and hence $\rho(\mathbf{R})\leq\theta\lambda_{0}$. Thus there exist $C>0$ and $\delta>0$ such that
$\|\mathbf{R}^{t}\|_{L^{1}\to L^{1}}\leq C(\lambda_{0}e^{-\delta})^{t}.$
Applying this to $P_{t}=\mathbf{K}^{t}P_{0}$ yields
$P_{t}=\lambda_{0}^{t}\mathbf{\Pi}P_{0}+\mathbf{R}^{t}P_{0}=\frac{\langle v_{0},P_{0}\rangle_{x}}{\langle v_{0},w_{0}\rangle_{x}}\lambda_{0}^{t}w_{0}+O(\lambda_{0}^{t}e^{-\delta t}),$
which is exactly (47). ∎
3.2. Reinterpretation of eigensystems by analogy with Markov chains
In understanding the sequence appearing in (40), the theory of Markov chains offers particularly valuable suggestions. In a Markov chain whose transition probability from state $j$ to state $i$ is denoted by $p_{ij}\geq 0$, the following quantity, called the taboo probability, is known (Chung, 1960).
(48) $p^{j}_{ij}(n):=$ $\mathbb{P}_{j}\left(X_{n}=i,\ X_{k}\neq j\ \text{for all}\ 1\leq k\leq n-1\right)$
(49) $=$ $\sum_{i_{1}\neq j}\sum_{i_{2}\neq j}\cdots\sum_{i_{k}\neq j}\cdots\sum_{i_{n-1}\neq j}p_{ii_{0}}p_{i_{0}i_{1}}\cdots p_{i_{k-1}i_{k}}\cdots p_{i_{1}j}$
It is well known that the following sequence constitutes the stationary distribution $\mu\left(i\right)$ of this Markov process.
(50) $\mu\left(i\right)=p_{ij}+\sum_{n=2}^{\infty}p^{j}_{ij}(n)$
If $\left(p_{ij}\right)_{1\leq i,j\leq M}$ is an irreducible stochastic matrix, the stationary distribution $\left(\mu\left(i\right)\right)_{1\leq i\leq M}$ can be equivalently described as the eigenvector corresponding to the largest eigenvalue 1. Considering the recurrence relation that the $n$-step taboo probabilities should satisfy, it can be formally expressed as follows:
(51) $p^{j}_{ij}(n)=\sum_{k=1}^{M}p_{ik}p^{j}_{kj}(n-1)-p_{ij}p^{j}_{jj}(n-1),\quad p^{j}_{ij}(1)=p_{ij}$
This relation does not require the matrix to be stochastic; an irreducible nonnegative matrix similarly yields an eigenvector corresponding to its Frobenius root (Oizumi et al., 2022a). Focusing on the right-hand side of (51), we see that the first term sums over all paths from every state $k$ to state $i$ at the previous step, while the second term subtracts the contribution of the paths that pass through state $j$. Then, by replacing the sum with an integral in (51), we observe that the resulting relation resembles the difference equation (42) that the coefficients of each power of $\lambda_{0}$ in (40) should satisfy.
However, from a measure-theoretic viewpoint, since a single point has Lebesgue measure zero, the expression (42) cannot be interpreted as “subtracting the paths passing through $y$ from all paths leading to xx.” Therefore, when $w(y)=1$, we define the series on the right-hand side of (40) as the direct contribution from $y$ to $x$. Similarly, we define the right-hand side of (43) as the adjoint direct contribution from $y$ to $x$. These two direct contributions respectively represent the degree of contribution from a past state $y$ to a future state $x$, and the degree of dependence of a future state $y$ on a past state $x$. Furthermore, we define the self-direct contribution as the direct contribution from a state $y$ to itself, where the direct contribution and its adjoint coincide. In a Markov process, a self-direct contribution of one indicates recurrence; in IPM, the value of $\lambda_{0}$ that makes the self-direct contribution equal to one gives the intrinsic growth rate.
(52) $\sum_{n=1}^{\infty}\frac{\Gamma_{n}\left(y_{0},y_{0}\right)}{\lambda_{0}^{n}}=\sum_{n=1}^{\infty}\frac{\Gamma_{n}^{*}\left(y_{0},y_{0}\right)}{\lambda_{0}^{n}}=1.$
In finite-dimensional models, namely transition matrix models, the self-direct contribution indeed reflects its name: it sums, over each number of steps, the paths that return to the same state for the first time.
3.3. Initial population dependence and expected contribution steps
Building on the previous section, the numerator $\langle v_{0},P_{0}\rangle_{x}$ of the expansion coefficient in (47)—the pairing of the reproductive value $v_{0}$ with the initial population $P_{0}$—quantifies the dependence of a future state $y$ on the initial distribution $P_{0}(\cdot)$:
(53) $\langle v_{0},P_{0}\rangle_{x}=c_{v}(y_{0})\sum_{n=1}^{\infty}\int_{\Omega^{d}}\frac{\Gamma_{n}^{*}(y_{0},x)\,P_{0}(x)}{\lambda_{0}^{\,n}}\,dx.$
To analyze $\langle v_{0},w_{0}\rangle$, the pairing of the reproductive value with the stable population distribution, we establish the following theorem.
Theorem 3.2.
Let $\Gamma_{m}^{*}(y_{0},x)$ satisfy (45) and $\Gamma_{n}(x,y_{0})$ satisfy (42). Then
(54) $\langle\Gamma_{m}^{*},\Gamma_{n}\rangle_{x}=\Gamma_{m+n}(y_{0},y_{0})+\Gamma_{m}(y_{0},y_{0})\,\Gamma_{n}(y_{0},y_{0}),\qquad m,n\geq 1.$
Proof.
To simplify the notation, we adopt the following expressions:
$\Gamma_{n}:=\Gamma_{n}(y_{0},y_{0})=\Gamma^{*}_{n}(y_{0},y_{0}),\qquad n\geq 1.$
Continuing, for integers $m\geq 1$ and $n\geq 1$ define
(55) $\phi(m,n):=\langle\Gamma_{m}^{*},\Gamma_{n}\rangle_{x}-\Gamma_{m+n}-\Gamma_{m}\,\Gamma_{n}.$
Using (42) and (45), we compute
$\langle\Gamma_{m}^{*},\Gamma_{n}\rangle_{x}$ $=\int_{\Omega^{d}}\left(\int_{\Omega^{d}}\Gamma_{m-1}^{*}(y_{0},\eta)\,K(\eta,x)\,d\eta-\Gamma_{m-1}\,K(y_{0},x)\right)\,\Gamma_{n}(x,y_{0})\,dx$
$=\langle\Gamma_{m-1}^{*},\Gamma_{n+1}\rangle_{x}+\int_{\Omega^{d}}\Gamma_{m-1}^{*}(y_{0},\eta)\,K(\eta,y_{0})\,\Gamma_{n}\,d\eta-\Gamma_{m-1}\int_{\Omega^{d}}K(y_{0},x)\,\Gamma_{n}(x,y_{0})\,dx$
(56) $=\langle\Gamma_{m-1}^{*},\Gamma_{n+1}\rangle_{x}+\Gamma_{m}\Gamma_{n}-\Gamma_{m-1}\Gamma_{n+1}.$
Substituting (56) into the definition (55) yields, for all integers $m\geq 1$ and $n\geq 1$, the diagonal-shift identity
$\phi(m,n)=\phi(m-1,n+1).$
Iterating this identity m1m-1 times gives $\phi(m,n)=\phi(1,m+n-1)$. But by direct computation with $m=1$ one verifies $\phi(1,k)=0$ for all $k\geq 1$, hence $\phi(m,n)=0$ for all $m,n\geq 1$. ∎
Remark 3.3.
The condition $y=x_{0}=y_{0}$ imposed in Theorem 3.2 is essential. If, instead, one keeps the assumptions of Theorem 2.6 and allows the variables in the direct contribution and its adjoint to vary independently, additional summation terms appear on the right-hand side of (54). Such terms not only obscure the biological interpretation of the eigensystem but also considerably complicate the associated computations. It is also worth noting that imposing $y=x_{0}=y_{0}$ alters the result at most by a multiplicative constant.
By invoking Theorem 3.2, $\langle v_{0},w_{0}\rangle$ is computed as follows:
(57) $\langle v_{0},w_{0}\rangle_{x}$ $=\int_{\Omega^{d}}v_{0}(y_{0},x)\,w_{0}(x,y_{0})\,dx$
(58) $=c_{v}(y_{0})\,c_{w}(y_{0})\int_{\Omega^{d}}\left(\sum_{m=1}^{\infty}\frac{\Gamma_{m}^{*}(y_{0},x)}{\lambda_{0}^{\,m}}\right)\left(\sum_{n=1}^{\infty}\frac{\Gamma_{n}(x,y_{0})}{\lambda_{0}^{\,n}}\right)\,dx$
(59) $=c_{v}(y_{0})\,c_{w}(y_{0})\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,m+n}}\int_{\Omega^{d}}\Gamma_{m}^{*}(y_{0},x)\,\Gamma_{n}(x,y_{0})\,dx$
[…]
Define the expected number of contributing steps by
(61) $\mathbb{E}_{y_{0}}[m]:=\sum_{n=1}^{\infty}n\,\frac{\Gamma_{n}(y_{0},y_{0})}{\lambda_{0}^{\,n}}.$
Then $(\ast)=\mathbb{E}_{y_{0}}[m]-\sum_{n\geq 1}\Gamma_{n}/\lambda_{0}^{n}=\mathbb{E}_{y_{0}}[m]-1$, hence (60) becomes
(62) $\langle v_{0},w_{0}\rangle_{x}=c_{v}(y_{0})\,c_{w}(y_{0})\,\mathbb{E}_{y_{0}}[m].$
Remark 3.4.
The quantity (61) admits the interpretation of an expected number of steps—under the probability distribution induced by the normalized self-contribution—required for past transitions to influence the current state. In the context of Markov processes, it corresponds to the expected return time. Since the kernel $K(\cdot)$ encodes state transitions (including birth and death processes), its specific interpretation is model dependent.
Substituting (40), (54), and (62) into (47) yields, in particular under the leading eigenvalue condition $\lambda_{0}=1$, demographic coefficients that determine the steady-state total population size:
(63) $\lim_{t\to+\infty}\int_{\Omega^{d}}P_{t}(x)\,dx=\frac{1}{\mathbb{E}_{y}[m]}\sum_{n=1}^{\infty}\sum_{k=1}^{\infty}\int_{\Omega^{d}}\Gamma_{k}(x,y_{0})\,dx\ \int_{\Omega^{d}}\Gamma_{n}^{*}(y_{0},\xi)\,P_{0}(\xi)\,d\xi.$
These coefficients represent the reproductive contribution of the initial population at state $y_{0}$, multiplied by the total direct contribution from state $y$ to all other states $x$, and normalized by the expected number of transition steps. Equivalently, the total population can be interpreted as the product of the expected reproduction and survival for the cohort at the initial state $y$ and the per-step contribution rate at state $y_{0}$. A larger expected step count $\mathbb{E}_{y_{0}}[m]$ implies a smaller per-step contribution of descendants. Since $\lambda_{0}=1$ corresponds to population replacement, (52) yields
(64) $\sum_{n=1}^{\infty}\Gamma_{n}(y_{0},y_{0})=1.$
We define
(65) $T_{y_{0}}:=\sum_{n=1}^{\infty}\Gamma_{n}(y_{0},y_{0}).$
Then the following proposition holds.
Proposition 3.5.
If $0<T_{y_{0}}\leq 1$, then $0<\lambda_{0}\leq 1$.
Proof.
This follows from the fact that the right-hand side of
$F(\lambda)=\sum_{n=1}^{\infty}\lambda^{-n}\Gamma_{n}(y_{0},y_{0})$
is strictly decreasing in $\lambda$ on and beyond the boundary of the spectral radius. ∎
Remark 3.6.
The quantity $T_{y_{0}}$ is referred to as the type reproduction number (TRN) (Heesterbeek and Roberts, 2007; Inaba, 2013) at state $y_{0}$. By analogy with Markov chains, it aggregates, over all $n$, the total contribution from individuals originating in state $y$ who either return to $y_{0}$ or produce descendants that reach $y_{0}$ for the first time at step $n$. From a measure-theoretic viewpoint, a single point in $\mathbb{R}^{d}$ has zero recurrence measure; thus this interpretation is merely heuristic. Nevertheless, given the meaning of the quantity and its relation to the dominant spectral value $\lambda_{0}$, it is natural to regard $T_{y_{0}}$ as the type reproduction number associated with a single state.
Thus far, analytical insight has been obtained through the spectral analysis of the discrete-time IPM (39), including the characteristic equation (52), the eigensystems (40) and (43), and the construction of the type reproduction number via Theorem 2.6. However, empirical IPMs often abstract away age-structured life history due to observational constraints, limiting biological interpretation. To address life-history, demographic, and evolutionary questions, it is therefore necessary to incorporate age structure explicitly into the mathematical formulation.
4. Multi-state McKendrick equation
4.1. Assumptions of Multi-state McKendrick equation
Multi-state age-structured IPM. We consider the transition kernel
$K:[0,\alpha)\times\Omega^{d}\times[0,\alpha)\times\Omega^{d}\to[0,\infty),\quad(a,x,s,y)\mapsto K(a,x\leftarrow s,y),$
representing the probability density of transitioning from state $y$ at age $s$ to state $x$ at age $a$. The kernel satisfies:
  • $K(a,x\leftarrow s,y)\in\mathbb{K}$ for $a>s$;
  • (Tail behavior) for all $N\in\mathbb{N}$ and $a>s$,
    (66) $\sup_{y\in\Omega^{d}}|x|^{N}K(a,x\leftarrow s,y)\to 0\quad\text{as }|x|\to\infty;$
  • $\lim_{a\downarrow s}K(a,x\leftarrow s,y)=\delta^{d}(x-y)$;
  • $K(a,x\leftarrow s,y)=0$ for $a<s$;
  • (Monotonicity) for $0\leq s\leq a_{0}\leq a_{1}<\alpha$,
    (67) $\int_{\Omega^{d}}K(a_{1},x\leftarrow s,y)\,dx\leq\int_{\Omega^{d}}K(a_{0},x\leftarrow s,y)\,dx;$
  • (Chapman–Kolmogorov equation) for $s\leq\tau\leq a<\alpha$,
    (68) $K(a,x\leftarrow s,y)=\int_{\Omega^{d}}K(a,x\leftarrow\tau,z)\,K(\tau,z\leftarrow s,y)\,dz;$
  • (Boundary) $K(\alpha,x\leftarrow s,y)=0$.
Note that $K(a,x\leftarrow s,y)$ is a (sub-)Markov transition density in the sense that
(69) $\int_{\Omega^{d}}K(a,x\leftarrow s,y)\,dx\leq 1.$
Let $P_{t}(a,x)$ denote the age-state density of a population at time $t\in[0,\infty)$, where $a\in[0,\alpha)$ is chronological age (with threshold $\alpha\in(0,\infty]$) and $x\in\Omega^{d}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional state variable. We define the multi-state age-structured IPM governed by $K$ by
(70) $P_{t+\varepsilon}(a+\varepsilon,x)=\int_{\Omega^{d}}K(a+\varepsilon,x\leftarrow a,y)\,P_{t}(a,y)\,dy,\quad\varepsilon>0,$
with initial condition
(71) $P_{0}(a,x)=\varphi(a,x)\in L^{1}\big([0,\alpha)\times\Omega^{d}\big),\quad\varphi(a,x)>0.$
Remark 4.1.
The monotonicity assumption in (67) is biologically motivated: it reflects the fact that cohort density decreases monotonically with age due to mortality.
To construct a renewal equation from (70), we introduce a fertility function $F:[0,\alpha)\times\Omega^{d}\times\Omega^{d}\to[0,\infty)$ that yields the inhomogeneous birth rate in age and state. Biologically, we assume:
  • $F(x\leftarrow y;a)>0$;
  • for each $a\in[0,\alpha)$, $F(x\leftarrow y;a)$ is measurable and continuous in $(x,y)\in\Omega^{d}\times\Omega^{d}$;
  • for each fixed $(a,y)$, $F(\cdot\leftarrow y;a)\in L^{1}(\Omega^{d})$;
  • if $\Omega^{d}$ is (partially) unbounded, let $\Omega^{\ell}\subseteq\Omega^{d}$ be an unbounded subset ($0\leq\ell\leq d$). There exist constants $C>0$, $\beta\in\mathbb{R}$ and $m_{i}\geq 0$ such that
    (72) $F(x\leftarrow y;a)\leq C\,\exp\{\beta a\}\left(1+\sum_{i=1}^{\ell}|y_{i}|^{m_{i}}\right),$
    where $y_{i}$ denotes the $i$th entry of $y\in\Omega^{\ell}$.
Once $F$ is defined, the generation of newborns is formulated by
(73) $P_{t}\left(0,x\right)=\int_{0}^{\alpha}\int_{\Omega^{d}}F(x\leftarrow y;a)\,P_{t}(a,y)\,dy\,da.$
Renewal equation. By the Chapman–Kolmogorov equation (68), (70) rewrites as
(74) $P_{t}(a,x)=\begin{cases}\int_{\Omega^{d}}K(a,x\leftarrow a-t,\xi)\,\varphi(a-t,\xi)\,d\xi,&\text{if }a\geq t,\\ \int_{\Omega^{d}}K(a,x\leftarrow 0,\xi)\,P_{t-a}(0,\xi)\,d\xi,&\text{if }a<t.\end{cases}$
Substituting (74) into (73) yields the renewal equation
(75) $P_{t}\left(0,x\right)$ $=G_{t}\left(x\right)+\int_{0}^{t}\mathbf{\Psi}\left(a\right)P_{t-a}\left(0,x\right)\,da,$
(76) $G_{t}\left(x\right)$ $:=\int_{t}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}F\left(x\leftarrow\xi;a\right)K\left(a,\xi\leftarrow a-t,y\right)\varphi\left(a-t,y\right)\,dy\,d\xi\,da,$
(77) $\mathbf{\Psi}\left(a\right)f\left(x\right)$ $:=\int_{\Omega^{d}}\int_{\Omega^{d}}F\left(x\leftarrow y;a\right)K\left(a,y\leftarrow 0,\xi\right)f\left(\xi\right)\,d\xi\,dy,$
for $f\in L^{1}\!\left(\Omega^{d}\right)$.
The Laplace transform is a standard tool for analyzing the asymptotics of renewal equations with integral operators (see (Inaba, 2017)). For $g\in L^{1}(0,\infty)$ and $r\in\mathbb{C}$, set
(78) $\hat{g}\left(r\right):=\int_{0}^{\infty}\exp\left\{-rt\right\}g\left(t\right)\,dt.$
Taking Laplace transforms in (75) gives
(79) $\hat{P}\left(0,x;r\right)=\hat{G}\left(x;r\right)+\hat{\mathbf{\Psi}}\left(r\right)\hat{P}\left(0,x;r\right).$
Here $\hat{\mathbf{\Psi}}(r):L^{1}(\Omega^{d})\to L^{1}(\Omega^{d})$ is
(80) $\hat{\mathbf{\Psi}}(r)f(x):=\int_{0}^{\alpha}\exp\{-ra\}\int_{\Omega^{d}}\int_{\Omega^{d}}F(x\leftarrow\xi;a)\,K(a,\xi\leftarrow 0,y)\,f(y)\,dy\,d\xi\,da.$
Formally solving (79) yields
(81) $\hat{P}(0,x;r)=\left(\mathbf{I}-\hat{\mathbf{\Psi}}(r)\right)^{-1}\hat{G}(x;r).$
Taking inverse Laplace transforms and applying Schumitzky and Wenska (1975) gives the asymptotic expansion (when $\psi(x,y;r)\in\mathbb{K}_{2}$,)
(82) $P_{t}(0,x)=G_{t}(x)+\sum_{k=0}^{\infty}\exp\{r_{k}t\}\frac{\langle v_{k}(0),\hat{G}(\cdot;r_{k})\rangle_{x}}{-\left\langle v_{k}(0),\left.\frac{d}{dr}\hat{\mathbf{\Psi}}(r)\right|_{r=r_{k}}w_{k}(0)\right\rangle_{x}}\,w_{k}(0,x).$
The parameters $r_{k}\in\mathbb{C}$ are ordered so that $\Re r_{k}\geq\Re r_{k+1}$ and are the singularities where the resolvent $\left(\mathbf{I}-\hat{\mathbf{\Psi}}(r)\right)^{-1}$ fails to exist.
The functions $v_{k}(0,\cdot)$ and $w_{k}(0,\cdot)$ denote the adjoint eigenfunction and the eigenfunction associated with $r_{k}$, respectively:
(83) $v_{k}(0,x)$ $=\int_{\Omega^{d}}v_{k}(0,\xi)\,\psi(\xi,x;r_{k})\,d\xi,$
(84) $w_{k}(0,x)$ $=\int_{\Omega^{d}}\psi(x,y;r_{k})\,w_{k}(0,y)\,dy,$
where
(85) $\psi(x,y;r):=\int_{0}^{\alpha}\exp\{-ra\}\int_{\Omega^{d}}F(x\leftarrow\xi;a)\,K(a,\xi\leftarrow 0,y)\,d\xi\,da.$
Remark 4.2.
The series in (82) generally diverges (Feller, 1941); in many cases, solutions cannot even be expressed by a series (Verduyn Lunel, 1989, 1990). However, if $\mathbf{\Psi}(a)$ is generated by a Hilbert–Schmidt kernel, the existence and uniqueness of solutions to (75) and the simplicity of the leading eigenvalue $r_{0}$ follow from Mode (1971).
4.2. The Dominant Root and Its Eigenstructure
We apply Theorem 2.6 to the eigenvalue system underlying (82).
Proposition 4.3.
Define
$\psi(x,y;r):=\int_{0}^{\alpha}\exp\left\{-ra\right\}\int_{\Omega^{d}}F(x\leftarrow\xi;a)\,K(a,\xi\leftarrow 0,y)\,d\xi\,da.$
Then $\psi\in\mathbb{K}$.
Proof of Proposition 4.3.
We verify that $\psi\in\mathbb{K}$.
Positivity. Since $F>0$ and $K\geq 0$, we have $\psi(x,y;r)>0$.
Integrability in $x$. Fix $y\in\Omega^{d}$ and choose $r>\beta$ so that $\exp\{-ra\}$ dominates the polynomial growth of $F$. By Tonelli,
$\int_{\Omega^{d}}\psi(x,y;r)\,dx$ $=\int_{0}^{\alpha}\exp\{-ra\}\int_{\Omega^{d}}\Bigl(\int_{\Omega^{d}}F(x\leftarrow\xi;a)\,dx\Bigr)\,K(a,\xi\leftarrow 0,y)\,d\xi\,da,$
which is finite since $F(\cdot\leftarrow\xi;a)\in L^{1}$ and $K\in\mathbb{K}$.
Essential boundedness in $y$. Fix $x\in\Omega^{d}$ and again take $r>\beta$. Then
$\psi(x,y;r)\leq\int_{0}^{\alpha}\exp\{-ra\}\int_{\Omega^{d}}F(x\leftarrow\xi;a)\,K(a,\xi\leftarrow 0,y)\,d\xi\,da,$
and the right-hand side is uniformly bounded in $y$ by the tail decay of $K$ and the $L^{1}$-bound on $F$.
Uniform $L^{1}$-boundedness in $y$. As above,
$\int_{\Omega^{d}}\psi(x,y;r)\,dx$ $\leq\int_{0}^{\alpha}\exp\{-ra\}\int_{\Omega^{d}}C_{F}(a,\xi)\,K(a,\xi\leftarrow 0,y)\,d\xi\,da,$
with $C_{F}(a,\xi):=Ce^{\beta a}(1+|\xi|^{M})$. Since $K$ decays faster than any polynomial in $\xi$, the integral is finite uniformly in $y$. Hence $\sup_{y}\int_{\Omega^{d}}\psi(x,y;r)\,dx<\infty$. ∎
Let $\psi^{(1)}:=\psi$ and, for $n\geq 1$,
$\psi^{(n+1)}(x,y;r_{0}):=\int_{\Omega^{d}}\psi(x,\xi;r_{0})\,\psi^{(n)}(\xi,y;r_{0})\,d\xi.$
Define, for $n\geq 1$,
$\psi_{n}(x,y;r_{0}):=\$$\psi^{(n)}(x,y;r_{0})$
(86) $+\sum_{\ell=1}^{n-1}(-1)^{\ell}\sum_{k=\ell}^{n-1}\psi^{(n-k)}(x,y;r_{0})\,\widehat{B}_{\,k,\,\ell}\!\Big(\psi^{(1)},\psi^{(2)},\dots,\psi^{(k)};y,y\Big)\Bigr|_{r=r_{0}}.$
Proposition 4.4.
Define $\Psi(r):=\sum_{n=1}^{\infty}\psi_{n}(y,y;r)$. Suppose there exists $r_{0}\in\mathbb{R}$ such that
(87) $\Psi(r_{0})=1.$
Then $r_{0}$ is a simple root of $\Psi(r)=1$, and there is no other root $r\in\mathbb{C}$ with $\Re r>r_{0}$. In particular, $r_{0}$ is the unique root with maximal real part (the dominant root).
Proof.
For $r$ with $0<\Psi(r)<1$, (38) gives
$\Psi(r)=\frac{((\mathbf{I}-\hat{\mathbf{\Psi}}(r))^{-1}\psi)(y,y;r)}{1+((\mathbf{I}-\hat{\mathbf{\Psi}}(r))^{-1}\psi)(y,y;r)}.$
By positivity, $\Psi$ is continuous and strictly decreasing in the real parameter $r$. Hence the equation $\Psi(r)=1$ has a unique real solution $r_{0}$, which is simple. If there existed a root $r_{k}\in\mathbb{C}$ with $\Re r_{k}>r_{0}$, then $\exp\{-\Re r_{k}a\}$ would yield a strictly smaller positive kernel than at $r_{0}$, so $\rho(\hat{\mathbf{\Psi}}(r_{k}))<\rho(\hat{\mathbf{\Psi}}(r_{0}))=1$, contradicting the singularity of $\mathbf{I}-\hat{\mathbf{\Psi}}(r_{k})$. Thus no such $r_{k}$ exists. ∎
From Theorem 2.6 and $\psi\in\mathbb{K}$, there exist $r_{0}>\beta$ and a nonzero, nonnegative function $w_{0}(x,y)\in L^{1}(\Omega^{d}\times\Omega^{d})$ such that
$w_{0}(x,y)=\int_{\Omega^{d}}\psi(x,\xi;r_{0})\,w_{0}(\xi,y)\,d\xi.$
Hence the stable multi-state age distribution at birth is
(88) $w_{0}(x,y)=w_{0}(y)\left(\sum_{n=1}^{\infty}\psi_{n}(x,y;r_{0})\right),\quad w_{0}(y)\neq 0.$
Similarly, there exist $r_{0}>\beta$ and a nonzero, nonnegative $v_{0}(y,x)\in L^{\infty}(\Omega^{d}\times\Omega^{d})$ such that
$v_{0}(y,x)=\int_{\Omega^{d}}v_{0}(y,\xi)\,\psi(\xi,x;r_{0})\,d\xi.$
Define, for $m\geq 1$,
$\psi_{m}^{*}(y,x;r_{0}):=\$$\psi^{(m)}(y,x;r_{0})$
(89) $+\sum_{\ell=1}^{m-1}(-1)^{\ell}\sum_{k=\ell}^{m-1}\widehat{B}_{\,k,\,\ell}\!\Big(\psi^{(1)},\psi^{(2)},\dots,\psi^{(k)};y,y\Big)\Bigr|_{r=r_{0}}\,\psi^{(m-k)}(y,x;r_{0}),$
and thus
(90) $v_{0}(y,x)=v_{0}(y)\left(\sum_{m=1}^{\infty}\psi_{m}^{*}(y,x;r_{0})\right),\quad v_{0}(y)\neq 0.$
4.3. Asymptotics and Demographic Interpretation
In the previous subsection we justified the dominant root $r_{0}$ and its eigenstructure. We now examine demographic indicators arising from this eigenstructure.
From (74) and (82), a Sharpe–Lotka–Feller-type asymptotic representation of the cohort density $P_{t}(a,x)$ is
$P_{t}(a,x)$ $=G_{t}(x)+\frac{\langle v_{0},\hat{G}\left(r_{0}\right)\rangle_{x}}{-\left\langle v_{0},\left.\frac{d}{dr}\hat{\mathbf{\Psi}}(r)\right|_{r=r_{0}}w_{0}\right\rangle_{x}}\exp\{r_{0}t\}\,w_{0}(a,x,y)$
(91) $\quad\times\left(1+O(\exp\{-\delta_{0}t\})\right),\quad\delta_{0}>0,$
where
$w_{0}(a,x,y):=\exp\{-r_{0}a\}\int_{\Omega^{d}}K(a,x\leftarrow 0,\xi)\,w_{0}(\xi,y)\,d\xi.$
Accordingly, (88) expresses the stable state distribution at age zero: $w_{0}(0,x,y)=w_{0}(x,y)$.
In the simple IPM (39), the direct contribution is a discrete sum over time steps. In the multistate McKendrick equation with continuous age, the kernel $\psi(x,y;r_{0})$ represents the total lifetime reproductive contribution of an individual starting from state $y$ to offspring with initial state $x$. The functions $\psi_{n}$ in (86), which constitute $w_{0}$, represent the contribution of each generation (index $n$).
Thus, in the model where $F(x\leftarrow\xi;a)$ determines both the number and the initial state of offspring, the stable density (88) aggregates contributions over all generations.
For the prefactor in (91), the quantity $\langle v_{0},\hat{G}\left(r_{0}\right)\rangle_{x}$ expands as
(92) $\langle v_{0},\hat{G}\left(r_{0}\right)\rangle_{x}$ $=\int_{\Omega^{d}}v_{0}(y,x)\,\hat{G}\left(x;r_{0}\right)\,dx$
$=\int_{\Omega^{d}}v_{0}(y,x)\int_{0}^{\alpha}\exp\{-r_{0}a\}\int_{a}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}$
(93) $\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow s-a,\eta)\,\varphi(s-a,\eta)\,d\eta\,d\xi\,ds\,da\,dx$
$=\int_{\Omega^{d}}v_{0}(y,x)\int_{0}^{\alpha}\int_{0}^{s}\exp\{-r_{0}a\}\int_{\Omega^{d}}\int_{\Omega^{d}}$
(94) $\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow s-a,\eta)\,\varphi(s-a,\eta)\,d\eta\,d\xi\,da\,ds\,dx$
$=\int_{0}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}v_{0}(y,x)\int_{0}^{s}\exp\left\{-r_{0}(s-\tau)\right\}\int_{\Omega^{d}}$
(95) $\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow\tau,\eta)\,\varphi(\tau,\eta)\,d\xi\,d\tau\,dx\,d\eta\,ds$
(96) $=\int_{0}^{\alpha}\int_{\Omega^{d}}v_{0}(\tau,y,\eta)\,\varphi(\tau,\eta)\,d\eta\,d\tau$
(97) $=:\langle v_{0},\varphi\rangle_{a,x},\quad(\tau,\eta)\to(a,x),$
where
(98) $v_{0}(a,y,x):=\int_{\Omega^{d}}v_{0}(y,\xi)\int_{a}^{\alpha}\exp\{-r_{0}(\tau-a)\}\int_{\Omega^{d}}F(\xi\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow a,x)\,d\eta\,d\tau\,d\xi.$
Finally, consider the denominator in (91):
$\left\langle v_{0},\frac{d}{dr}\hat{\mathbf{\Psi}}(r)\big|_{r=r_{0}}w_{0}\right\rangle_{x}.$
Using the representations of $v_{0}$ and $w_{0}$ at age zero,
$-\left\langle v_{0},\frac{d}{dr}\hat{\mathbf{\Psi}}(r)\big|_{r=r_{0}}w_{0}\right\rangle_{x}=-\int_{\Omega^{d}}\int_{\Omega^{d}}v_{0}\left(y,x\right)\,\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\,w_{0}\left(\xi,y\right)\,d\xi\,dx$
(99) $=-v_{0}\left(y\right)w_{0}\left(y\right)\sum_{n=1}^{\infty}\sum_{m=1}^{n}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}\left(y,x;r_{0}\right)\,\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\,\psi_{m}\left(\xi,y;r_{0}\right)\,d\xi\,dx,$
where
(100) $\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}=\int_{0}^{\alpha}a\,\exp\left\{-r_{0}a\right\}\int_{\Omega^{d}}F(x\leftarrow\eta;a)\,K(a,\eta\leftarrow 0,\xi)\,d\eta\,da.$
To normalize, set
$\langle v_{0},\hat{\mathbf{\Psi}}\left(r_{0}\right)w_{0}\rangle_{x}$ $=\langle v_{0},w_{0}\rangle_{x}$
(101) $=v_{0}\left(y\right)w_{0}\left(y\right)\,E_{n}\left(y\right),$
where the mean contributing generation number is
(102) $E_{n}\left(y\right):=\sum_{n=1}^{\infty}n\,\psi_{n}\left(y,y;r_{0}\right).$
Dividing (99) by (102) yields the average generation interval
$\bar{g}_{\mathrm{L}}:=\frac{-1}{E_{n}\left(y\right)}\sum_{n=1}^{\infty}\sum_{m=1}^{n}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}\left(y,x;r_{0}\right)\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\psi_{m}(\xi,y;r_{0})\,d\xi\,dx.$
Using the interpretations of (98) and (90), equation (97) yields the direct contribution $\bar{R}_{\mathrm{L}}(y)$ from the initial distribution $\varphi$ to the descendants of $y$:
$\bar{R}_{\mathrm{L}}(y):=\frac{\langle v_{0},\varphi\rangle_{a,x}}{v_{0}\left(y\right)}.$
We may also define cohort-based quantities in terms of the spectral radius $\lambda_{0}:=\Lambda\big(\hat{\mathbf{\Psi}}(0)\big)$ of the next-generation operator $\hat{\mathbf{\Psi}}(0)$:
(103) $E_{n}^{c}(y)$ $:=\sum_{n=1}^{\infty}\frac{n}{\lambda_{0}^{n}}\,\psi_{n}(y,y;0),$
(104) $\bar{g}_{0}$ $:=\frac{1}{E_{n}^{c}(y)}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\sum_{m=1}^{n}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}(y,x;0)\,\left.\frac{d}{dr}\psi(x,\xi;r)\right|_{r=0}\,\psi_{m}(\xi,y;0)\,d\xi\,dx,$
$\bar{R}_{0}(y)$ $:=\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\int_{0}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n}^{*}(y,\xi;0)$
(105) $\quad\times\int_{a}^{\alpha}\int_{\Omega^{d}}F(\xi\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow a,x)\,\varphi(a,x)\,d\eta\,d\tau\,d\xi\,dx\,da.$
In this multistate setting, $\lambda_{0}$ corresponds to the basic (net) reproduction number (Inaba, 2017). The average life expectancy adjusted for the population growth rate is
$e_{0}\left(r_{0}\right):=\frac{\int_{0}^{\alpha}\int_{\Omega^{d}}w_{0}\left(a,x,y\right)\,dx\,da}{\int_{\Omega^{d}}w_{0}\left(0,\xi,y\right)\,d\xi}.$
As a new demographic indicator, we define the per-generation total contribution of the cohort with initial state $y$, denoted by $\Upsilon(y,r_{0})$, as follows.
(106) $\Upsilon(y,r_{0}):=\frac{\int_{\Omega^{d}}w_{0}\left(0,\xi,y\right)\,d\xi}{w_{0}\left(y\right)E_{n}\left(y\right)}.$
At replacement level ($r_{0}=0$ so $\lambda_{0}=1$),
$E_{n}\left(y\right)=E_{n}^{c}\left(y\right),\quad\bar{g}_{\mathrm{L}}=\bar{g}_{0},\quad\bar{R}_{\mathrm{L}}\left(y\right)=\bar{R}_{0}\left(y\right).$
Thus, at replacement level, the stationary population is characterized by the contribution to descendants with a specific initial state $y$, the generation interval, the life expectancy at birth, and the total contribution of the cohort with initial state $y$, consistent with the classical McKendrick/Leslie theory. Furthermore, the introduction of the average contributory generation number $E_{n}^{c}(y)$ together with $\Upsilon(y,0)$ provides genealogical resolution beyond earlier models.
4.4. Other Demographic Indicators and Statistical Quantities
Similarly, the type reproduction number is the direct contribution from an ancestor with the same initial condition $y$:
(107) $T_{y}=\sum_{n=1}^{\infty}\psi_{n}(y,y;0).$
Corollary 4.5.
By the definition of the basic reproduction number $\lambda_{0}$,
(108) $1=\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\psi_{n}(y,y;0).$
Hence, if $\lambda_{0}=1$, then $T_{y}=1$ and $r_{0}=0$. Moreover, by the monotonicity of $\left(1-\frac{1}{\lambda}\hat{\mathbf{\Psi}}(0)\right)^{-1}$ in $\lambda$, it follows that $\lambda_{0}<1$ implies $T_{y}<1$ and $r_{0}<0$.
Associated probabilities (definitions). We recall the two associated probability measures on $0,\alpha)$:
(109) $\mathbb{P}_{\mathrm{L}}([0,a);y)$ $=\frac{1}{E_{n}(y)}\sum_{n=1}^{\infty}\sum_{m=1}^{n}\int_{0}^{a}\exp\{-r_{0}\tau\}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi^{*}_{\,n-m}(y,x;r_{0})$
$\qquad\times\int_{\Omega^{d}}F(x\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow 0,\xi)\,\psi_{m}(\xi,y;r_{0})\,d\eta\,d\xi\,dx\,d\tau,$
(110) $\mathbb{P}_{0}([0,a);y)$ $=\frac{1}{E_{n}^{c}(y)}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,n}}\sum_{m=1}^{n}\int_{0}^{a}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi^{*}_{\,n-m}(y,x;0)$
$\qquad\times\int_{\Omega^{d}}F(x\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow 0,\xi)\,\psi_{m}(\xi,y;0)\,d\eta\,d\xi\,dx\,d\tau,$
with the normalizing constants
$E_{n}(y):=\sum_{n=1}^{\infty}n\,\psi_{n}(y,y;r_{0}),\qquad E_{n}^{c}(y):=\sum_{n=1}^{\infty}\frac{n}{\lambda_{0}^{\,n}}\,\psi_{n}(y,y;0).$
Cumulant expansions under the associated probabilities. Write the expectations under (109)–(110) as
$\mathbb{E}_{y}[\cdot]:=\int(\cdot)\,\mathbb{P}_{\mathrm{L}}(da;y),\qquad\mathbb{E}_{y}^{c}[\cdot]:=\int(\cdot)\,\mathbb{P}_{0}(da;y).$
Define the cumulant generating functions
$\Theta_{\mathrm{L}}(t;y):=\ln\mathbb{E}_{y}\!\left[\exp\{ta\}\right],\qquad\Theta_{0}(t;y):=\ln\mathbb{E}_{y}^{c}\!\left[\exp\{ta\}\right],$
and the cumulants $\kappa_{k}(y):=\partial_{t}^{k}\Theta_{\mathrm{L}}(0;y)$, $\kappa_{k}^{c}:=\partial_{t}^{k}\Theta_{0}(0;y)$ $(k\geq 1)$. Then, for $r$ near $r_{0}$ and $r$ near 0, respectively,
(111) $\ln\!\Biggl(\frac{\langle v_{0},\,\hat{\mathbf{\Psi}}(r)\,w_{0}\rangle_{x}}{v(y)\,w(y)}\Biggr)$ $=\ln E_{n}(y)\;+\;\Theta_{\mathrm{L}}\!\big(-\,(r-r_{0};y)\big)$
$=\ln E_{n}(y)\;+\;\sum_{k=1}^{\infty}\frac{\kappa_{k}}{k!}\,(-1)^{k}\,(r-r_{0})^{k},$
(112) $\ln\!\Biggl(\frac{\big\langle\bar{v}_{0},\,\frac{1}{\lambda_{0}}\hat{\mathbf{\Psi}}(r)\,\bar{w}_{0}\big\rangle_{x}}{\bar{v}(y)\,\bar{w}(y)}\Biggr)$ $=\ln E_{n}^{c}(y)\;+\;\Theta_{0}\!\big(-\,r;y\big)$
$=\ln E_{n}^{c}(y)\;+\;\sum_{k=1}^{\infty}\frac{\kappa_{k}^{c}}{k!}\,(-1)^{k}\,r^{k}.$
Second-order truncation (generation-time statistics). Retaining only the first two cumulants in (111)–(112) yields
(113) $\ln\!\Biggl(\frac{\langle v_{0},\,\hat{\mathbf{\Psi}}(r)\,w_{0}\rangle_{x}}{v(y)\,w(y)}\Biggr)=\ln E_{n}(y)\;-\;\bar{g}_{\mathrm{L}}\,(r-r_{0})\;+\;\frac{1}{2}\,\sigma_{\mathrm{L}}^{2}\,(r-r_{0})^{2}\;+\;o\!\big((r-r_{0})^{2}\big),$
(114) $\ln\!\Biggl(\frac{\big\langle\bar{v}_{0},\,\frac{1}{\lambda_{0}}\hat{\mathbf{\Psi}}(r)\,\bar{w}_{0}\big\rangle_{x}}{\bar{v}(y)\,\bar{w}(y)}\Biggr)=\ln E_{n}^{c}(y)\;-\;\bar{g}_{0}\,r\;+\;\frac{1}{2}\,\sigma_{0}^{2}\,r^{2}\;+\;o\!\big(r^{2}\big),$
where the generation-time mean and variance under the associated probabilities are
$\bar{g}_{\mathrm{L}}$ $:=\kappa_{1}=\mathbb{E}[a],\quad\sigma_{\mathrm{L}}^{2}:=\kappa_{2}=\mathbb{V}_{y}(a),$
$\bar{g}_{0}$ $:=\kappa_{1}^{c}=\mathbb{E}^{c}[a],\quad\sigma_{0}^{2}:=\kappa_{2}^{c}=\mathbb{V}^{c}(a).$
Thus, the linear term encodes the (cohort- or lineage-based) mean generation time, and the quadratic term encodes its variance.
Remark 4.6.
Accordingly, any representative demographic indicator derived from the multi-state McKendrick equation inherently reflects the entire sequence of intergenerational transitions and cannot be characterized solely by cohort-based quantities, as in the classical McKendrick or Leslie models.
5. Discussion
This paper develops a determinant-free way to describe the dominant eigenstructure of positive integral projection models (IPMs) through a reference-point construction. The central outcome is that, in the dominant spectral regime, solutions of the Fredholm equation admit an explicit representation as series of iterated kernels organized by a fixed reference pair $(x_{0},y_{0})$. This provides a continuous-state analogue of Euler–Lotka-type characteristic equations while keeping the interpretation close to what matters in applications: how population-scale growth and stable structure are shaped by the accumulation of genealogical contributions across generations.
From discrete taboo probabilities to a continuous-state analogue.
A key conceptual issue in transferring “taboo” ideas from discrete models to continuous ones is that avoiding a single point is a null event under Lebesgue measure, and pointwise taboo events are therefore not meaningful if one works only with $L^{1}$-type objects. In contrast, in discrete-time matrix models, paths that avoid a designated state form a nontrivial family, and this makes taboo probabilities a natural tool for decomposing eigenvectors at the dominant growth rate. The role of taboo probabilities is thus best understood as a path decomposition principle: split genealogical paths by whether they hit a designated state first, and reorganize the renewal structure that results.
The reference-point construction in this paper implements an analogous decomposition directly at the kernel level. Rather than viewing taboo as a literal event in a continuous state space, we use a rank-one correction built from point evaluation at $(x_{0},y_{0})$ and show that the resulting recursions unify the discrete-time IPM setting and the continuous-time multi-state McKendrick setting under a common linear architecture in mathematical biology. The detailed analogy with taboo probabilities (as used in Markov chains) and the meaning of this “continuousization” are discussed in the application sections, where the construction can be read in terms of genealogical paths and first-visit decompositions.
Interpretation in simple IPMs: stable structure as genealogical aggregation.
In a simple discrete-time IPM, the dominant eigenvalue and its eigenfunctions are often introduced as abstract objects controlling asymptotic growth and stable trait distributions. The determinant-free representation developed here makes these objects more concrete: the stable distribution and the reproductive value are expressed through iterated kernels that can be read as multi-step life-history contributions. This viewpoint aligns with how IPMs are used in ecology and demography: one is typically interested in which states and which sequences of transitions contribute most to long-run growth, and how perturbations of survival, development, or reproduction at particular states propagate through the system. Because the expansion is organized by a reference point, it naturally supports decompositions “conditioned on visiting” a focal state and highlights how specific states act as gateways for genealogical flow.
Interpretation in multi-state McKendrick equations: demographic indicators driven by ancestry and initial state.
The multi-state McKendrick equation provides a continuous-time, age-structured framework in which state variables beyond age (e.g., type, location, stage, or other traits) interact with aging, survival, reproduction, and transitions. In such settings, demographic quantities are inherently genealogical: the population size and composition at time $t$ are determined not merely by current cohort properties but by how ancestral histories and state-switching trajectories accumulate over generations.
Within this perspective, the present framework emphasizes that indicators such as expected births, generation time, and related life-history summaries are not simply “cohort-wise” quantities. They depend on the distribution of ancestral pathways that feed into present states, especially when state transitions are not uniquely determined by age (for example, models with migration, switching among health states, or transitions among social or economic categories). By making the influence of iterated genealogical contributions explicit, the reference-point moments yield quantities that directly quantify how strongly the current demographic regime is shaped by the ancestral past. In particular, indicators such as $E_{n}(y)$ summarize the average magnitude of generational contributions inherited from ancestry, and they are naturally interpreted as measures of “how many generations back” remain demographically influential for individuals currently in state $y$.
Why a determinant-free formulation matters in applications.
Classical determinant-based Fredholm formulations originating in Fredholm (1903) are conceptually fundamental, but they can be difficult to compute and do not directly reflect the path-based reasoning that practitioners use when interpreting structured population models. Applied work therefore often turns to quadrature-based matrix approximations (Ellner and Rees, 2006; White et al., 2016), which can be effective but require care to maintain qualitative features of the continuous model. In contrast, the present approach replaces determinant expressions by explicit sums over iterated kernels within the biologically natural kernel class $\mathbb{K}$. This shifts the emphasis from determinants to genealogical aggregation and supports interpretation in terms of life-history trajectories and generation-by-generation contributions.
Limitations and future directions (biological emphasis).
Several directions remain open. First, we have focused on the dominant spectral value and its eigenfunctions, which are the quantities most directly tied to long-run growth and stable structure; extending the discussion to non-dominant spectral components would be relevant for transient dynamics, cohort resonance, and responses to temporal variability. Second, identifying families of kernels for which the reference-point moments (and hence the resulting demographic indicators) can be computed or approximated efficiently is important for empirical deployment, especially in models with multiple interacting states.
Finally, because the framework makes the influence of all past generations explicit, it suggests a natural way to formulate and compare life-history strategies in terms of how they redistribute genealogical contributions across states and generations. In evolutionary ecology, this may support new ways to analyze selection on state-dependent reproduction and transition schedules. More broadly, whenever dominant spectral quantities are used as measures of growth, persistence, or “fitness,” the present genealogical expansions may offer an interpretation in which long-run outcomes are traced back to concrete ancestral pathways rather than to determinant expressions.
Acknowledgements
I would like to express my sincere gratitude to Kumiko Oizumi, Shin Oizumi, Ko Oizumi, and Hiroko Oizumi for their warm support and encouragement throughout this research. I also wish to thank Hisashi Inaba for his valuable advice, and Youichi Enatsu for the insightful discussions that greatly contributed to this study. We would further like to pay tribute to the late Professor Nobuhiko Fuji, who passed away in December 2025, and to express our heartfelt gratitude for the many fruitful discussions and the wealth of mathematical knowledge he shared with us.
Funding
This work was supported by a Health and Labour Sciences Research Grant, titled ”Research contributing to the perspective for demographic and social change in the post COVID-19 era” (Grant Number 23AA2005). Yuki Chino acknowledge the support by NSTC grant: 111-2115-M-A49-009-MY3.